### v1

# A Survey of World Models for Autonomous Driving

Abstract
Recent breakthroughs in autonomous driving have revolutionized the way vehicles perceive and interact with their surroundings. In particular, world models have emerged as a linchpin technology, offering high-fidelity representations of the driving environment that integrate multi-sensor data, semantic cues, and temporal dynamics. Such models unify perception, prediction, and planning, thereby enabling autonomous systems to make rapid, informed decisions under complex and often unpredictable conditions. Research trends span diverse areas, including 4D occupancy prediction and generative data synthesis, all of which bolster scene understanding and trajectory forecasting. Notably, recent works exploit large-scale pretraining and advanced self-supervised learning to scale up models’ capacity for rare-event simulation and real-time interaction. In addressing key challenges – ranging from domain adaptation and long-tail anomaly detection to multimodal fusion – these world models pave the way for more robust, reliable, and adaptable autonomous driving solutions. This survey systematically reviews the state of the art, categorizing techniques by their focus on future prediction, behavior planning, and the interaction between the two. We also identify potential directions for future research, emphasizing holistic integration, improved computational efficiency, and advanced simulation. Our comprehensive analysis underscores the transformative role of world models in driving next-generation autonomous systems toward safer and more equitable mobility.
Index Terms:
Autonomous Driving, World Models, Self-Supervised Learning, Behavior Planning, Generative Approaches
1 Introduction
1.1 Overview
The quest for fully autonomous driving has rapidly become a global focal point in both scientific research and industry endeavors. At its core lies the ambition to simultaneously reduce traffic accidents, alleviate congestion, and enhance mobility for diverse societal groups [1]. Current statistics underscore that human error remains the principal cause of accidents on the road [2], indicating that minimizing human intervention could significantly lower the incidence of traffic-related fatalities and injuries. Beyond safety, economic factors (e.g., reducing congestion and optimizing logistics) further propel the development of autonomous driving technologies [3].
Despite these compelling incentives, achieving high-level autonomy demands overcoming substantial technical hurdles. Foremost among these is perceiving and understanding dynamic traffic scenarios, which requires fusing heterogeneous sensor streams (e.g., LiDAR, radar, cameras) into a cohesive environmental representation [4, 5]. From complex urban layouts to high-speed highways, autonomous vehicles must rapidly assimilate multi-modal data, detect salient objects (vehicles, pedestrians, cyclists), and anticipate their motion under varying conditions – such as inclement weather, unstructured roads, or heavy traffic [6, 7]. Furthermore, real-time decision-making introduces stringent computational constraints, imposing millisecond-level responsiveness to address unexpected obstacles or anomalous behaviors in the driving environment [8, 9]. Equally pivotal is the system’s resilience in extreme or long-tail scenarios (e.g., severe weather, construction zones, or erratic driving behaviors), where performance shortfalls can compromise overall safety [10, 11].
Within this context, constructing robust and stable world models has emerged as a foundational element. The notion of a world model involves creating a high-fidelity representation of the driving environment – encompassing static structures (e.g., roads, buildings) and dynamic entities (e.g., vehicles, pedestrians) [3, 8]. A comprehensive world model continuously captures semantic and geometric information while updating these representations in real-time, thereby informing downstream tasks such as physical world prediction [12, 13]. Recent advances integrate multi-sensor data to refine these representations, such as generative approaches [14, 15] that simulate the physical world for training that unify heterogeneous sensor inputs into consistent top-down perspectives [16, 17].
In turn, these robust world models leverage environmental representations to optimize the behavior planning of intelligent agents, providing the keystone for safer and more efficient autonomous driving applications. By enabling proactive trajectory optimization, real-time hazard detection, and adaptive route planning, they directly mitigate risks posed by unforeseen hazards [5] and align with evolving vehicle-to-everything (V2X) systems [9]. Ultimately, world models facilitate more cohesive integration between perception and control subsystems, streamlining the closed-loop autonomy pipeline [18, 19].
Existing surveys on world models that involve autonomous driving can generally be classified into two categories. The mainstream category focuses on describing general world models that find applications across multiple fields [20, 21, 22], with autonomous driving being just one of the specific areas. The second category [23, 24], concentrates on the application of world models within the autonomous driving sector, and attempts to summarize the current state of the field. There are only a few existing surveys on world models in autonomous driving, they tend to broadly categorize these studies and often focus solely on world simulation or lack discussions on the interaction between behavior planning and physical world prediction, resulting in a lack of a clear taxonomy in the field. In this paper, we aim not only to define and categorize world models for autonomous driving formally but also to provide a comprehensive review of recent technical progress and explore their extensive applications in various sectors, particularly emphasizing their transformative potential in autonomous driving. This structured taxonomy allows us to highlight how these models are shaped by and adapt to the challenges of the automotive sector.
1.2 Contributions
Guided by the principle that the world model is central to the understanding of dynamic scenes, this survey aims to provide a comprehensive, structured review of existing methodologies. We categorize state-of-the-art research into three key areas: Future Prediction of the Physical World: Focusing on the physical world evolution of both dynamic objects and static entities [11, 25]; Behavior Planning for Intelligent Agents: Examining generative and rule-based planning methods that produce safe, efficient paths under uncertain driving conditions [12, 13]; Interaction Between Behavior Planning and Future Prediction: Highlighting how unified frameworks can capture agent interactions and leverage predictive insights for collaborative optimization [26, 27, 18]. Specifically, we provide:
  • An In-Depth Analysis of Future Prediction Models: We discuss how Image-/BEV-/OG-/PC-based methods achieve geometric and semantic fidelity in dynamic scenes, including 4D occupancy forecasting and diffusion-based generation.
  • Investigation of Behavior Planning: We explore the behavior planning through both rule-based and learning-based approaches, demonstrating notable improvements in robustness and collision avoidance.
  • Proposition of Interactive Model Research: We systematically review interactive models that jointly address future prediction and agent behavior, indicating how this synergy can vastly enhance real-world adaptability and operational safety.
We conclude by identifying open challenges, such as seamless integration of self-supervised approaches [26], large-scale simulation for rare-event augmentation [10, 28], and real-time multi-agent coordination [27], offering directions for future exploration. With the expanding research landscape and the urgency of real-world adoption, this survey aspires to serve as a valuable reference point for researchers and practitioners, laying the groundwork for safer, more robust autonomous driving solutions.
§1§2§3§4§5§6§7§8§3.1§3.2§3.3§4
Figure 1: Structure of the overall review. The first row shows the structure of the paper. The second and third rows detail future prediction of the physical world, behavior planning for intelligent agents, and interaction between behavior planning and future prediction. The fourth row highlights various methodologies for training models in autonomous driving, covering both self-supervised learning paradigms, pretraining strategies, and innovative approaches for data generation.
1.3 Structure
A summary of the structure of this paper can be found in Fig. 1, which is presented as follows: Sec. 1 introduces the significance of world models in autonomous driving and outlines the societal and technical challenges they address. Sec. 2 provides a comprehensive background on the formulation and core tasks of world models in autonomous driving, specifically focusing on the future prediction of the physical world and behavior planning for intelligent agents. Sec. 3 details the taxonomy of methods: Sec. 3.1 delves into methods for future prediction of the physical world, discussing physical world evolution of dynamic objects and static entities. Sec. 3.2 discusses advanced behavior planning approaches that emphasize the generation of safe, effective driving strategies. Sec. 3.3 investigates the interactive relationship between future prediction and behavior planning, highlighting collaborative optimization techniques for complex scenarios. Sec. 4 explores different approaches to data and training paradigms, including supervised and self-supervised learning, and data generation techniques. Sec. 5 examines the application areas and tasks where world models can be applied, discussing the impact of these technologies across diverse domains including perception, prediction, simulation, and system integration. Sec. 6 provides a detailed evaluation of world models for autonomous driving, assessing their effectiveness across various tasks and metrics. Sec. 7 explores open challenges, potential research avenues, and promising directions for further innovation in autonomous driving technologies. Sec. 8 concludes the survey and summarizes key findings, reiterating the importance of robust world models for autonomous driving.
2 Background
In this section, we first provide a detailed problem formulation for world models in autonomous driving (Sec. 2.1), encompassing two key tasks: future prediction of the physical world and behavior planning for intelligent agents. Then, in Sec. 2.2, we introduce key terminologies and concepts relevant to world models, such as representation spaces, generative models, and spatiotemporal modeling techniques. These aspects lay the foundation for understanding state-of-the-art methods.
2.1 Problem Formulation
2.1.1 Core Tasks in World Models
In autonomous driving, a critical aspect is accurately predicting the future states of both the ego vehicle and its surrounding environment. To address the core tasks, world models $\bm{w}$ in autonomous driving takes sensor inputs (including a set of multi-view images $\bm{I}$ and a set of LiDAR points $\bm{P}$) collected from previous frames and infers the scene and trajectory for the next frames. Specifically, the ego trajectory at time $T+1$, denoted as $\tau^{T+1}$, is predicted alongside the surrounding scene $\bm{z}^{T+1}$. $\bm{w}$ models the coupled dynamics of the ego vehicle’s motion and the environment’s evolution. Formally, the function $\bm{w}$ is given by:
$\!\!\!\!\bm{z}^{T+1},\!\tau^{T+1}\!=\!\bm{w}((\bm{I}^{T},\!\cdots,\!\bm{I}^{T-t}),\!(\bm{P}^{T},\!\cdots,\!\bm{P}^{T-t})).$ (1)
The first core task is future prediction of the physical world [29, 27, 17, 30], which involves forecasting the future states of dynamic entities such as vehicles, pedestrians, and traffic elements. This task emphasizes capturing potential interactions, stochastic behaviors, and uncertainties within rapidly changing and complex scenes. Advanced techniques, such as 4D occupancy prediction and generative models, play a crucial role in addressing this challenge by leveraging multi-modal sensor data and probabilistic forecasting frameworks. The second core task is behavior planning for intelligent agents [18, 31, 19, 32], focusing on generating optimal and feasible trajectories for the ego vehicle. It requires accounting for safety constraints, dynamic obstacles, traffic regulations, and real-time adaptability. Behavior planning is often achieved through a combination of model-based and learning-based approaches that ensure robustness and responsiveness in diverse driving scenarios.
2.2 Context and Terminology
2.2.1 Representation Spaces
Occupancy Grid (OG) Representation. An OG representation partitions the environment into discrete cells, each annotated with a probability of occupancy, thereby offering a unified representation for static and dynamic objects in 3D space. Although OG approaches are highly descriptive, they typically require large memory and computational resources, which can limit their applicability in real-time autonomous systems [29].
Bird’s-Eye View (BEV) Representation. A BEV Representation converts multi-modal sensor data into a top-down view, facilitating more intuitive spatial understanding, particularly for motion prediction and trajectory planning. However, BEV representations may have difficulty capturing fine-grained 3D geometries, especially in environments with complex depth relationships [27, 14].
Point Cloud (PC) Representation. A PC Representation uses raw 3D point data collected from LiDAR sensors to encode the spatial and geometric structure of the environment. Point clouds provide fine-grained 3D details and are inherently suited for capturing both static and dynamic objects in high-resolution environments. Despite their precision, point cloud processing is computationally intensive due to data sparsity and the high dimensionality of the input [17].
2.2.2 Generative Models
Generative models, such as VAEs and diffusion architectures [33, 14], play a pivotal role in simulating future driving environments by facilitating trajectory prediction, rare-event synthesis, and uncertainty modeling through diverse scenario generation. For instance, OccSora [34] introduces a diffusion-based 4D occupancy generation model that yields realistic, temporally consistent 3D driving simulations, while InfinityDrive [28] pushes temporal limits by producing long-duration, high-resolution video sequences of future states. Despite these advancements, balancing high-fidelity outputs with computational efficiency remains an active research challenge.
2.2.3 Spatiotemporal Modeling Techniques
Spatiotemporal modeling addresses how entities evolve in space and time within dynamic driving environments.
Transformers. employ attention mechanisms to capture long-range temporal dependencies and multi-agent interactions, facilitating improved motion prediction and scene understanding [35].
Slot Attention, on the other hand, segments the environment into interpretable entities, enabling more structured representations for downstream tasks such as behavior planning and motion forecasting [11]. These approaches have proven valuable in addressing dynamic and complex driving scenarios. However, developing models that are both computationally efficient and robust in real-time settings remains a core challenge.
This section provides a foundational understanding of world models in autonomous driving, highlighting how representation spaces, generative methods, and advanced modeling techniques together address the critical issues of uncertainty, complexity, and real-time operation.
3 Taxonomy
In this section, we classify the methodologies presented in the uploaded works into three primary categories, highlighting their unique contributions and interconnections: (1) Future Prediction of the Physical World, (2) Behavior Planning for Intelligent Agents, and (3) Interaction Between Behavior Planning and Future Prediction. This section provides a structured understanding of the diverse approaches in autonomous driving research. Future prediction methods focus on modeling the future states of static structures and dynamic entities, behavior planning optimizes vehicle trajectories, and their integration bridges the gap between environmental understanding and actionable decision-making. These frameworks collectively address the multifaceted challenges of autonomous navigation, driving innovation in safety, efficiency, and robustness.
3.1 Future Prediction of the Physical World
3.1.1 Image Representations
Image representation-based approaches to future prediction in autonomous driving capitalize on advanced generative models to address challenges like limited data availability and complex scene dynamics. Techniques such as diffusion models and generative adversarial networks (GANs) enable the synthesis of high-fidelity images, enriching training datasets and enhancing the robustness of downstream perception and planning modules.
For instance, DrivingDiffusion [14] deploys a diffusion model to generate temporally coherent video frames, creating diverse training scenarios that bolster object detection and motion segmentation in dynamic environments. By employing cascaded generation modules and localized prompts, DrivingDiffusion maintains spatial and temporal consistency, which is pivotal for reliably modeling traffic agents in motion. DriveDreamer [11] further integrates image synthesis and trajectory prediction to simulate realistic urban driving conditions, especially high-density intersections. This focus on interactive elements and unpredictable traffic flows enriches the planning process, enabling vehicles to anticipate and adapt to sudden changes on the road. Meanwhile, ViDAR [17] leverages GANs to synthesize future frames from monocular video sequences, employing temporal consistency layers to mitigate artifacts in moving objects and adapt to varying environmental factors such as weather changes.
Building on these foundations, DrivingWorld [36] proposes a GPT-based video generation framework that unifies detection, prediction, and planning tasks within a single model, thereby streamlining data processing and decision-making. Complementarily, HoloDrive [37] synthesizes 2D–3D multi-modal scenes to reinforce the alignment between image and LiDAR data, enhancing predictive accuracy for more robust reconstruction of complex driving environments. ReconDreamer [38] focuses on the real-time reconstruction of dynamic scenes, emphasizing online restoration techniques that enable responsive modeling of evolving traffic conditions. In simulation environments, Imagine-2-Drive [39] advances high-fidelity world modeling in CARLA, equipping researchers and practitioners with more realistic test scenarios to evaluate autonomous systems under diverse conditions. Finally, Delphi [40] moves toward end-to-end autonomous driving by generating controllable, long-duration videos that facilitate broader generalization in complex traffic scenarios.
Collectively, these image-based generative frameworks significantly strengthen the adaptability and safety of autonomous vehicles. They not only supply synthetic but highly realistic training data for perception tasks, but also refine predictive capabilities and planning strategies, enabling intelligent vehicles to navigate with greater awareness and robustness in rapidly shifting driving contexts.
3.1.2 BEV Representations
Bird’s-Eye View (BEV) representations have emerged as a powerful paradigm for modeling future states in autonomous driving environments, thanks to their ability to unify multi-modal sensor data and facilitate structured, top-down scene understanding. As an alternative or complement to image-based approaches, BEV methods reduce redundant perceptual details from raw sensor readings, rendering high-level, spatially coherent views that are well-suited for motion prediction, occupancy modeling, and trajectory forecasting.
A prominent illustration of BEV’s capabilities is GenAD [19], which introduces an end-to-end generative motion prediction pipeline via tokenized scene representations and a Variational Autoencoder (VAE). By capturing interactions between the ego vehicle and surrounding agents in a structured BEV space, GenAD yields accurate trajectory predictions even in cluttered traffic scenarios. Extending this representation to account for evolving semantic occupancy, BEVWorld [27] integrates a diffusion model to predict the geometric and semantic evolution of forthcoming scenes, merging multi-modal inputs (e.g., LiDAR and images) into a continuous latent space. In parallel, UNO [29] tackles 4D occupancy prediction from a purely unsupervised standpoint, leveraging historical signals to generate future point clouds and occupancy grids in a continuous fashion. Its implicit decoder mitigates quantization pitfalls and achieves state-of-the-art outcomes on benchmarks such as Argoverse and nuScenes. Similarly, FIERY [30] exploits spatiotemporal convolutional networks atop BEV inputs to model the dynamic evolution of traffic entities; by integrating probabilistic forecasting, FIERY can output a spectrum of possible future scenarios, which is particularly valuable for safety-critical tasks like instance segmentation and motion prediction.
In addition to these representative works, a host of recent studies further highlights the diversity and adaptability of BEV-based world models. Methods like BevDepth [41] and PowerBEV [42] refine depth estimation and instance prediction within the BEV framework, whereas BevControl [43] and BEVGen [44] push the boundaries of multi-perspective consistency and generative street-view rendering, respectively. At the same time, InfinityDrive [28], GaussianWorld [45], and NeMo [46] expand upon volumetric, generative, and streaming occupancy strategies, illustrating the versatility of BEV-centric approaches when combined with next-generation neural architectures. Collectively, these advances underscore how BEV representations not only offer a unified canvas for perception, prediction, and planning but also significantly improve robustness in dynamic driving environments, paving the way for safer and more reliable autonomous systems.
Generative scene synthesis has recently emerged as a valuable tool for enriching the data landscape used in behavior planning. For instance, BEVGen [44] synthesizes multi-view images from BEV segmentation masks, thereby enhancing data augmentation for tasks such as 3D detection and tracking. By manipulating the bird’s-eye view layout, BEVGen offers a structured way to generate diverse driving scenarios – helping learning-based planners acquire greater robustness and adaptability.
Similarly, BEVControl [43] focuses on simplifying scene editing while maintaining geometric and visual consistency in the top-down view. This tool provides intuitive visual guidance for planning by allowing practitioners to modify complex elements of the scene with minimal effort, thus accelerating the iterative design and testing of learned planning strategies.
Generative scene synthesis also contributes to behavior planning by enriching the data and visual references available. BEVGen [44] synthesizes multi-view images from BEV segmentation masks, enhancing data augmentation for tasks like 3D detection and tracking. BEVControl [43] simplifies scene editing while ensuring geometric and visual consistency, providing intuitive visual guidance for planning. These learning-based approaches underscore the transformative potential of data-driven behavior planning frameworks in addressing the complexities of autonomous driving, enabling safer and more adaptive decision-making in dynamic environments.
3.1.3 OG Representations
Occupancy grid (OG) representations have become a cornerstone for predicting future states in autonomous driving, primarily due to their capacity to encode high-fidelity 3D spatial information. By discretizing the environment into voxel grids, these methods offer a more detailed and geometrically consistent view than projections like BEV, making them especially well-suited for modeling scene evolution and capturing fine-grained interactions over time. However, such high-resolution modeling often incurs substantial memory and computational costs, especially in large-scale or real-time scenarios. The potential for sparse sensor data or irregular updates further complicates prediction accuracy, and controlled scenario generation typically relies on carefully curated trajectory conditions.
A foundational approach in this domain is OccWorld [47], which introduces 4D occupancy grids to jointly predict semantic occupancy and agent trajectories in cluttered driving environments. Building on this groundwork, MUVO [48] integrates multi-modal data (i.e., LiDAR point clouds and camera streams) into high-resolution 3D voxel grids, elevating prediction quality in both vision-based and point cloud tasks. Extending temporal modeling, OccSora [34] leverages diffusion models to generate future occupancy states, enabling controlled scenario simulation for motion-planning pipelines. Further integrating prediction with decision-making, Drive-OccWorld [49] fuses occupancy forecasts with a motion planner, leveraging historical grid states and vehicle trajectories to optimize driving strategies in real-world contexts.
Recent works push OG-based methods toward greater efficiency and scalability. DFIT-OccWorld [50] employs a decoupled dynamic flow mechanism and image-based supervision, achieving end-to-end, non-autoregressive occupancy forecasting while mitigating computational overhead. GaussianWorld [45] reframes OG prediction in a 4D Gaussian space, explicitly modeling scene evolution to enhance streaming performance on datasets like nuScenes without substantial resource demands. Across these advances, learning-based OG techniques consistently prove their worth in motion forecasting, scene reconstruction, and safe route planning. By balancing high-resolution occupancy outputs with emerging strategies for real-time efficiency, the latest research points toward solutions that can seamlessly adapt to the complexity and dynamism of real-world autonomous driving.
3.1.4 PC Representations
3D point cloud (PC) representations, typically obtained from LiDAR sensors, have emerged as a linchpin in autonomous driving for tasks such as occupancy forecasting, dynamic scene modeling, and predictive reconstruction. Their ability to capture detailed geometric layouts of vehicles, pedestrians, and surrounding infrastructure has proven indispensable. However, the sparsity and irregular sampling of LiDAR scans – combined with real-time computational constraints – continue to pose significant algorithmic challenges.
Early generative efforts targeted realistic LiDAR data simulation for data augmentation and testing. For instance, lidarGeneration [51] adopted a GAN-based framework to mimic real-world LiDAR statistics, while Lidarsim [7] introduced LidarSim to construct near-realistic LiDAR scenes by leveraging empirical sensor distributions. Subsequent works further advanced the diversity and efficiency of point cloud generation. Lidargen [52] proposed a deep learning pipeline that preserves physical consistency while increasing point cloud variability. Meanwhile, AXform [53] employed attention mechanisms to map latent features into 3D space, and PSF [54] developed a straight-flow approach that accelerates the generation process without compromising fidelity. In parallel, NFL [55] introduced Neural LiDAR Fields for synthesizing novel-view point clouds, demonstrating effectiveness under dynamic or partially observed scenes. Extending neural rendering concepts, Nerf-lidar [56] proposed NeRF-LiDAR, leveraging neural radiance fields to produce high-resolution LiDAR data suitable for realistic 3D reconstructions.
A foundational step toward future prediction with PC data is offered by 4DOcc [57], who recast point cloud forecasting as a 4D occupancy prediction problem. Rather than directly inferring raw point coordinates, they map sequential LiDAR scans into intermediate occupancy grids and employ temporal models (e.g., Transformers) to estimate future occupancy states. This paradigm neatly bypasses heavy reliance on meticulously annotated LiDAR data by rendering pseudo LiDAR sequences from the predicted occupancy volumes.
Recent work further demonstrates the synergy of point cloud representations with advanced neural volumetric techniques. For instance, NeMo [46] leverage self-supervised 3D reconstruction and motion cues to achieve robust geometric and dynamic understanding – often outperforming 2D or BEV-based approaches in scenarios with frequent occlusions or varied terrains. Similarly, ViDAR [17] extends this idea by predicting future point cloud distributions from purely visual inputs, thus unifying semantic, spatial, and temporal learning. By integrating generative modeling with latent-space transformations, ViDAR demonstrates notable gains in 3D detection, motion forecasting, and route planning.
Collectively, these methodologies highlight the effectiveness of 3D point cloud representations in autonomous driving. By enabling precise modeling of spatial geometry and dynamic interactions, they address core challenges in scene understanding, motion prediction, and safety-critical evaluation – paving the way for scalable, resilient autonomous systems capable of navigating increasingly complex real-world environments.
3.2 Behavior Planning for Intelligent Agents
Behavior planning for intelligent agents aims to generate safe and effective driving strategies and trajectory plans based on the current environmental state and predicted dynamics. As the ultimate goal, a motion planner needs to plan a safe and comfortable trajectory towards the target point. The general idea for motion planners is to output the most likely trajectory given a sampling of possible candidates and semantic results from preceding modules [58, 59, 60, 61]. This process involves leveraging high-level decisions and environmental understanding to ensure that the chosen trajectory meets safety and comfort requirements while adapting to dynamic conditions.
For implicit methods [62, 63, 59], the network directly generates planned trajectories or control commands. Although such designs are direct and simple, they suffer from robustness issues and a lack of interpretability. In light of an interpretable spirit, explicit methods usually build a cost map with a trajectory sampler to generate the desired trajectory by choosing the optimal candidate with the lowest cost. For instance, cost volume based planners [64, 65, 58] score and rank future ego-vehicle trajectories by constructing cost volumes that reflect the confidence [64, 65, 58] in different trajectory options using a specific form of trajectory modelling within a sampler, which allows the planner to evaluate and select the most promising trajectory based on predicted cost metrics and thus identify the optimal path that balances safety, efficiency, and adherence to driving objectives.
3.2.1 Rule-Based Planning
Rule-based behavior planning is guided by predefined heuristics and algorithms, which offer interpretability, facilitate debugging, and ensure predictable decision-making. Although these approaches often perform well in structured or relatively stable driving conditions, they can encounter challenges when faced with significant uncertainty or rapidly changing scene dynamics. Rule-based planners have been extensively explored in the literature [66, 67, 68, 69, 70] and are widely adopted due to their safety guarantees and transparency [71, 72, 73, 74, 75]. Leveraging the current position, velocity, and distance to the lead vehicle, rule-based planners compute longitudinal acceleration to progress safely toward a target. IDM [76] represents a classic, non-learning-based vehicle motion planning algorithm that uses graph-based search to reach the target and a PID velocity controller to avoid collisions. Dauner et al. [77] refine IDM by sampling multiple trajectories and rolling out a constant velocity world model to choose the trajectory with the lowest cost.
Conventional trajectory optimization methods generally aim to determine a complete path from the initial configuration to the final goal. However, given the inherently dynamic and uncertain nature of real-world driving environments, accurate long-horizon plans cannot be devised beforehand. Consequently, model-predictive control (MPC) has emerged as a prominent technique for real-time path planning [78, 79, 80, 81, 82] by iteratively minimizing a cost function and selecting a locally optimal trajectory at each timestep.
Motion planning is often formulated as an optimization problem that minimizes a hand-engineered cost function to produce an optimal trajectory [83, 84, 85, 86]. To streamline this process, some approaches either assume a quadratic objective or split the planning task into lateral and longitudinal components. Methods like A* [87], RRT [88], and dynamic programming [85] commonly search for optimal solutions. CoverNet [89] generates a set of trajectories, uses cost functions to evaluate them, and selects the lowest-cost option. Although these techniques stand out for their parallelizability, interpretability, and guaranteed functionality, they can be less robust in real-world environments and often demand significant hyperparameter tuning. In cost volume-based planning, the cost map may be crafted manually [65, 58, 90] using intermediate representations such as segmentation outputs or HD maps, or it can be learned directly from the network [64]. DSDNet [60] merges handcrafted and learned cost components to construct a unified cost volume. ST-P3 [91] follows a similar principle by combining both approaches to select the most promising trajectory and identify the likeliest candidate, assisted by high-level commands and without HD maps. The cost function makes full use of the learned occupancy probability field (i.e., segmentation maps in Prediction) and other pre-existing knowledge, such as traffic rules, to guarantee the final trajectory’s safety and smoothness.
More recently, PFBD [12] has integrated planning features into a deep reinforcement learning framework, aiming to synchronize higher-level decision-making with lower-level trajectory planning. While the underlying policy network is trained via reinforcement learning, rule-based constraints derived from path-planning topologies help maintain consistency between decision and planning layers. By using rule-based topological trajectories as a foundation and learning an optimal policy through deep reinforcement learning, PFBD effectively incorporates learning-based methods into the autonomous vehicle’s core modules of behavioral decision making and trajectory planning.
3.2.2 Learning-Based Planning
Autonomous driving, particularly in urban environments, demands that vehicles interact with a wide array of traffic participants [92, 93, 94] and navigate through intricate and ever-changing traffic conditions. Traditional rule-based planning methods, which rely heavily on manually-designed heuristics, often fall short in addressing such challenges due to their inability to account for the exhaustive range of edge-case scenarios [95, 96]. Moreover, as the decision framework grows more complex, ensuring compatibility between newly added and existing rules becomes increasingly problematic [97]. In response to these limitations, learning-based planners have emerged as a promising alternative, leveraging data and computational power to scale towards fully autonomous driving. By utilizing data-driven models like deep neural networks, reinforcement learning algorithms, and large language models, these planners are better equipped to handle uncertainty and complexity in dynamic traffic environments. While they often outperform rule-based methods in terms of adaptability and scalability, they also pose new challenges in interpretability, generalization, and safety assurance. The following explores learning-based planning, highlighting key contributions.
A central concern in learning-based approaches is balancing performance with safety. SafeDreamer [98] integrates Lagrangian methods into the Dreamer framework to form a safe reinforcement learning algorithm, achieving near-zero safety violations across multiple tasks in the Safety-Gymnasium benchmark. This highlights a growing trend toward incorporating explicit safety constraints into RL-based planners. Another reinforcement learning–based approach, Think2Drive [99], introduces a neural planner capable of handling quasi-realistic scenarios using techniques such as the reset technique, automated scenario generation, termination-priority replay strategy, and steering cost function. This approach demonstrates that an agent can “think” through imagined scenarios in a latent world model to enhance learning efficiency. Tested in CARLA-v2, Think2Drive achieves higher sample efficiency and improved decision quality compared to standard RL baselines.
A parallel thread focuses on large language models. DrivingGPT [100] unifies world modeling and planning under a multimodal autoregressive Transformer, framing driving decisions as a next-token prediction task. This design outperforms strong baselines on both video generation and end-to-end planning in nuPlan and NAVSIM. DRIVESIM [101] investigates how large language models, when augmented with vision or sensor data, can serve as internal world models for driving tasks. Itegrating multimoda large language models into autonomous vehicles can enhance vehicle intelligence and user interaction by leveraging real-time data (e.g., traffic, weather) to improve awareness and navigation. They facilitate user-friendly communication for planning and personalize driving settings. However, experimental study reveals that while they excel at interpreting individual images, they struggle to synthesize coherent narratives across frames, leading to considerable inaccuracies in understanding trajectory planning.
In cost volume-based planning, cost maps can be constructed using learning-based methods [58, 102] to represent the confidence levels of trajectories within a sampler’s specific trajectory modeling framework. DSDNet [60] integrates both hand-crafted and learning-based costs to form a comprehensive cost volume. Similarly, ST-P3 [91] employs this combination to select the optimal trajectory. The cost function leverages the learned occupancy probability field (segmentation maps in prediction) and extensive prior knowledge to ensure the safety and smoothness of the final trajectory. Drive-OccWorld [49] introduces an occupancy-based cost function, where the learned-volume cost is inspired by ST-P3 [91]. It utilizes a learnable head based on learned bird’s-eye view representations to generate a cost volume, providing a more comprehensive evaluation of the complex environment.
As state-of-the-art systems integrate increasingly rich sensor data and more powerful generative or language models, their capacity to handle edge cases, covariate shifts, and diverse traffic conditions continues to expand. Yet, questions of interpretability and large-scale deployment persist, underscoring the ongoing need for both innovative algorithms and robust real-world validations.
3.3 Interaction Between Behavior Planning and Future Prediction
The integration of behavior planning and future prediction plays a pivotal role in enhancing decision-making efficiency and safety in complex, dynamic scenarios. By coupling predictive insights (e.g., how other agents or the environment might evolve) with an autonomous vehicle’s own actions, these methods emphasize the complementary relationship between accurately modeling the physical world’s future states and generating intelligent, context-aware behaviors.
Early foundational tools like CARLA [103] and provide essential simulation platforms to support research on both perception and planning. While these works primarily focus on environment modeling, they lay the groundwork for experiments in interaction-aware planning, in which predictive models of other road users inform the ego-vehicle’s trajectory. Later research introduces more policy-oriented perspectives [1], underscoring the societal and infrastructural considerations critical to deploying planning and prediction frameworks at scale.
Recent advancements have propelled generative and predictive models to the forefront of integrated behavior planning and future prediction paradigms. BEVWorld [27] introduces a novel approach by tokenizing multimodal sensor inputs into a unified and compact Bird’s Eye View (BEV) latent space for environment modeling. This model employs a multi-modal tokenizer to encode information and a latent BEV sequence diffusion model to predict future scenarios, effectively integrating perception, prediction, and planning tasks. Similarly, PowerBEV [42] presents a powerful yet lightweight framework for instance prediction in BEV. By utilizing a parallel, multi-scale module built from lightweight 2D convolutional networks, PowerBEV predicts future instances in a spatio-temporally consistent manner, informing kinematic models about impending scenarios. Complementing these approaches, DriveDreamer [11] pioneers a world model entirely derived from real-world driving scenarios. It leverages a two-stage training pipeline to understand structured traffic constraints and anticipate future states, enhancing closed-loop performance by synchronizing future prediction modules with strategy optimization. Extensive experiments on the nuScenes benchmark validate DriveDreamer’s capability in generating precise, controllable driving videos and realistic driving policies, thereby advancing the integration of behavior planning and future prediction in autonomous driving systems.
Building upon foundational work, recent approaches have increasingly integrated prediction and planning in autonomous driving systems. For instance, DriveWorld [18] introduces a 4D pre-trained scene understanding model that processes multi-camera driving videos in a spatiotemporal manner. This model employs a Memory State-Space Model, comprising a Dynamic Memory Bank for learning temporal-aware latent dynamics and a Static Scene Propagation module for spatial-aware latent statics, effectively unifying BEV representations for perception, prediction, and path planning. Similarly, Drive-OccWorld [49] proposes a vision-centric 4D forecasting world model tailored for end-to-end planning in autonomous driving. It introduces semantic and motion-conditional normalization within its memory module to accumulate semantic and dynamic information from historical BEV embeddings. These features are then utilized by the world decoder to forecast future occupancy and flow, directly influencing the ego vehicle’s trajectory decisions. These advancements underscore the growing trend of integrating prediction and planning, enhancing the adaptability and safety of autonomous driving systems.
4 Data and Training Paradigms
This section focuses on the methodologies for training models in autonomous driving, emphasizing self-supervised learning paradigms, pretraining strategies, and innovative approaches for data generation.
4.1 Self-Supervised Learning for World Models
Early work on self-supervised learning for world models in autonomous driving focuses on pure spatio-temporal prediction [104], which estimates future LiDAR scans without explicit tracking. Building on that, UnO [29] and EO [105] introduce occupancy-based representations: UnO uses 4D occupancy fields to model dynamic changes, while EO learns future occupancy by comparing predicted and actual LiDAR sweeps. RenderWorld [106] then integrates multi-view images via a Gaussian-based Img2Occ module to self-supervise 3D occupancy labels, pushing occupancy modeling further by unifying 2D and 3D cues. Subsequently, UniPAD [26] and ViDAR [17] elevate pre-training paradigms: UniPAD harnesses volumetric differentiable rendering for robust 2D–3D representation learning, and ViDAR aligns temporal visual features with point clouds to improve scene understanding and motion forecasting. In parallel, COPILOT4D [107] embraces a discrete diffusion approach to generate future states from tokenized sensor data, capturing complex distributions without labeled samples. For end-to-end driving, SSR [108] introduces a sparse scene representation that leverages only a handful of navigation tokens with a self-supervised temporal alignment module, while CarFormer [35] utilizes slot attention in a BEV framework to learn object-centric representations. This end-to-end trend is further advanced by the LAW [109], which predicts future latent features conditioned on ego actions to reduce supervision needs. Finally, BEVWorld [27] unifies diverse sensor modalities within a single BEV latent space using diffusion modeling, showcasing the current push toward holistic, self-supervised world models that fuse complementary signals for perception, prediction, and planning in autonomous driving.
4.2 Pretraining Strategies
Large-Scale Pretraining Frameworks. Large-scale pretraining has emerged as a powerful mechanism for boosting robustness and generalization in autonomous driving systems. By learning rich semantic and geometric representations from vast amounts of data, these approaches offer significant advantages prior to any domain-specific fine-tuning. Notably, ViDAR++ [110] fuses multi-modal sensor data (e.g., LiDAR, cameras) with high-level semantic cues, underlining the value of large-scale pretraining in real-world environments. Meanwhile, UniPAD [26] bridges 2D and 3D representations through differentiable voxel rendering, achieving superior performance across diverse sensor modalities compared to conventional 3D self-supervised methods. Extending this predictive paradigm, UniWorld [111] leverages pre-trained world models to anticipate future states, thereby improving adaptability in dynamic settings and enhancing scene understanding, even under sparse or noisy conditions.
Building on these, BEVWorld [27] integrates multi-modal inputs into a cohesive latent space, highlighting the synergy between semantic and geometric features and demonstrating robust generalization when transitioning from simulation to real-world scenarios. Additionally, DriveWorld [18] employs a 4D scene understanding framework that pre-trains on large-scale multi-camera driving videos, effectively bridging the gap between simulated and real-world environments to bolster reliability under edge-case conditions.
Collectively, these large-scale pretraining frameworks illustrate a clear developmental progression: from early predictive models combining multi-modal data, to universal paradigms that unify voxel rendering and 2D appearance-based methods. By providing rich, pre-trained representations of driving scenes, they reduce the need for extensive manual annotations and improve performance on downstream tasks. As a result, large-scale pretraining paves the way for safer, more robust, and more generalizable autonomous driving systems.
4.3 Data Generation for Training
Data generation is essential for creating diverse and realistic datasets to enhance model training.
Generative Models for Data Synthesis. Generative models rapidly become an essential tool for data synthesis in autonomous driving, creating diverse, high-fidelity scenarios that bolster downstream tasks such as perception, prediction, and control. Recent works highlight multi-stage or conditional pipelines to enhance realism and expand coverage of edge-case conditions. For instance, MagicDrive3D [112] leverages controllable 3D generation and any-view rendering to handle geometrically complex street scenes, while OccSora [34] introduces diffusion-based 4D occupancy generation for long-sequence, trajectory-aware simulations. Other approaches (e.g., Panacea [1], ReconDreamer [38]) focus on panoramic or high-quality video generation to ensure temporal coherence in complex driving maneuvers.
Simulation-conditioned methods like SimGen [113] integrate real-world and simulator data to enhance scene diversity, whereas large-scale GPT-style models (e.g., DrivingWorld [36]) unify video generation and planning in a single paradigm. High-fidelity 2D–3D generation further emerges in HoloDrive [37] and Imagine-2-Drive [39], which incorporate diffusion-based policy actors for robust trajectory planning. Pushing temporal limits, InfinityDrive [28] extends generation to multi-minute horizons, enabling comprehensive scenario “storylines” (e.g., progressive traffic buildup).
Beyond these, DrivingDojo [114] provides richly annotated data for interactive, action-conditioned video generation, while OOD-centric strategies (e.g., OODGen [115]) address safety-critical out-of-distribution scenarios. Large Language Models (LLMs) also play an expanding role: DriveDreamer-2 [116] and DriveDreamer4D [25] integrate textual prompts to interpret user-defined maneuvers, boosting scenario diversity and training robustness. Parallel research explores controllable occupancy diffusion (DOME [117]) and multi-view volume-aware processes (WoVoGen [118]), each contributing to realism, consistency, and adaptability. Evaluations like WORLDSIMBENCH [119] assess perceptual and control-level metrics for generated sequences, promoting more rigorous standards in data-driven simulations. Collectively, these generative methods advance the synthesis of complex driving environments – ranging from congested urban settings to extreme weather – significantly enhancing the adaptability and safety of autonomous driving systems.
5 Application Areas and Tasks
Through the integration of world models, autonomous driving systems have made significant progress in critical tasks such as scene understanding, motion prediction, simulation, and end-to-end driving, demonstrating greater reliability and adaptability.
5.1 Scene Understanding
Scene understanding serves as the foundation for real-time environmental perception and comprehension in autonomous driving. This technology enables autonomous systems to identify roads, vehicles, pedestrians, and traffic signs, thereby facilitating safe and effective driving decisions. With the advancement of the field, world models are introduced to enhance the depth and breadth of scene understanding. By integrating data from various sensors, world models build 3D representations of the environment and predict future scene evolution, bridging information gaps while improving safety and efficiency.
By fusing information from multiple sensors such as cameras and LiDAR, world models reconstruct the 3D structure of the environment, enabling accurate recognition and localization of roads, vehicles, and pedestrians. For example, ViDAR [17] and GaussianWorld [45] integrate multi-modal sensor data into a unified 3D scene to achieve precise semantic segmentation and reliable object detection under various road layouts. Moreover, world models simulate the continuous evolution of 3D scenes to enhance the prediction of dynamic environmental changes. For instance, OccWorld [47] improves occupancy forecasting accuracy by capturing the temporal progression of 3D scenarios, thereby boosting downstream tasks such as collision avoidance and path planning. In summary, world models offer powerful tools for scene understanding in autonomous driving by leveraging multi-modal data fusion, continuous scene evolution modeling, and occupancy-based generative frameworks. This integration significantly enhances perception, prediction, and decision-making capabilities in complex driving environments.
5.2 Motion Prediction
In autonomous driving, motion prediction is pivotal for forecasting the future trajectories of surrounding entities, such as vehicles and pedestrians. Accurate predictions enable autonomous systems to make informed decisions, ensuring safe and efficient navigation through dynamic environments. Traditional approaches often rely on modular pipelines that handle perception, prediction, and planning tasks separately. However, these methods can suffer from accumulated errors and limited inter-module communication, potentially hindering adaptability to novel scenarios. To address these challenges, recent advancements have explored the integration of world models – comprehensive representations of the environment that encapsulate both spatial and temporal dynamics – into motion prediction frameworks. One notable example is TrafficBots [120], a multi-agent policy framework that formulates data-driven traffic simulation as a world model. By introducing navigational information and time-invariant latent personality traits for each agent, TrafficBots can simulate realistic multi-agent behaviors, enhancing the planning capabilities of autonomous vehicles. Another approach is OccWorld [47], which employs a 3D semantic occupancy representation to model the development of driving scenes. OccWorld predicts the evolution of both dynamic agents and static elements, facilitating tasks like 4D occupancy forecasting and trajectory planning without relying on extensive human-annotated labels. These methodologies demonstrate the potential of world models to enhance motion prediction by providing a unified, holistic understanding of the driving environment, leading to more robust and adaptable autonomous systems.
5.3 Simulation
Simulation plays a pivotal role in autonomous driving by providing controlled environments that can test diverse scenarios without real-world risks. However, traditional simulation methods often fail to capture the full complexity of dynamic driving conditions, prompting researchers to explore world models – comprehensive frameworks that encompass both spatial and temporal aspects of the driving context. For instance, the OG Representation-based OccSora [34] is a diffusion-based 4D occupancy generation model designed to simulate the evolution of 3D scenes for autonomous driving. OccSora employs a 4D scene tokenizer to extract compact spatio-temporal representations, thereby enabling high-fidelity reconstruction of extended occupancy videos. By learning a diffusion transformer on these representations, OccSora generates 16-second videos with realistic 3D layouts and temporal consistency, demonstrating an understanding of the spatial and temporal distributions in driving scenes. This trajectory-aware 4D generation serves as a world simulator to inform decision-making in autonomous vehicles. Another method, image Representation-based SimGen [113] learns to produce diverse driving scenes by integrating data from both simulators and the real world. Through a cascade diffusion pipeline, SimGen bridges the gap between simulated and real-world data by adhering to simulator-driven layout guidance and rich text prompts, ultimately yielding realistic driving scenarios. By combining simulated and real-world data, SimGen boosts the diversity and authenticity of generated scenes, which is essential for training robust autonomous driving systems. These advancements in world model-driven simulation highlight the potential for creating more realistic and adaptable virtual environments, thereby strengthening the robustness and reliability of autonomous driving solutions.
5.4 End-to-End Driving
End-to-end driving refers to systems that directly map raw sensor inputs to driving actions using deep learning models, bypassing traditional modular pipelines. This approach aims to streamline the decision-making process, potentially improving reaction times and reducing error propagation inherent in segmented systems. However, challenges such as data efficiency, interpretability, and adaptability to diverse driving scenarios persist. To address these issues, recent advancements have integrated world models – comprehensive representations of the driving environment – into end-to-end frameworks.
For instance, the NMP [64] introduces a model that not only predicts driving actions but also provides interpretable intermediate representations, enhancing transparency in decision-making processes. Similarly, SSR [108] investigates the necessity of explicit perception modules within end-to-end systems, contributing to the discourse on model architecture optimization. These approaches demonstrate that incorporating world models into end-to-end driving frameworks can enhance performance, robustness, and interpretability, marking a significant step toward more reliable autonomous vehicles.
TABLE I: 3D Occupancy prediction performance on the Occ3D-nuScenes [121, 122] validation set (Sec. 6.2).
  Methods GT mIoU $\uparrow$
Others
barrier
bicycle
bus
car
cons. veh
motorcycle
pedestrian
traffic cone
trailer
truck
dri. sur
other flat
sidewalk
terrain
manmade
vegetation
TPVFormer [123] 3D 27.83 7.22 38.90 13.67 40.78 45.90 17.23 19.99 18.85 14.30 26.69 34.17 55.65 35.47 37.55 30.70 19.40 16.78
BEVFormer [124] 3D 26.88 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46
OccFormer [125] 3D 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97
CTF-Occ [121] 3D 28.53 8.09 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.0
[…]
OccSora [34] 3D 27.4 11.7 22.6 0.0 34.6 29.0 16.6 8.7 11.5 3.5 20.1 29.0 61.3 38.7 36.5 31.1 12.0 18.4
OccLLaMA [126] 3D 75.2 65.0 87.4 93.5 77.3 75.1 60.8 90.7 88.6 91.6 67.3 73.3 81.1 88.9 74.7 71.9 48.8 42.4
DOME [117] 3D 83.1 36.6 90.9 95.9 85.8 92.0 69.1 95.3 96.8 92.5 77.5 86.8 93.6 94.2 89.0 85.5 72.2 58.7
[…]
RenderOcc [127] 2D 23.93 5.69 27.56 14.36 19.91 20.56 11.96 12.42 12.14 14.34 20.81 18.94 68.85 33.35 42.01 43.94 17.36 22.61
SurroundOcc [128] 2D 20.30 - 20.59 11.68 28.06 30.86 10.70 15.14 14.09 12.06 14.38 22.26 37.29 23.70 24.49 22.77 14.89 21.86
GaussianFormer [129] 2D 19.10 - 19.52 11.26 26.11 29.78 10.47 13.83 12.58 8.67 12.74 21.57 39.63 23.28 24.46 22.99 9.59 19.12
GaussianOcc [130] 2D 9.94 - 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 - 20.10 17.58 8.61 10.29
OccNeRF [131] 2D 9.53 - 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 - 20.81 24.75 18.45 13.19
SelfOcc [132] 2D 9.30 0.00 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 0.00 26.30 26.54 14.22 5.60
RenderWorld [106] 2D 27.87 6.83 32.54 7.44 21.15 29.92 16.68 11.43 17.45 16.48 24.02 27.86 75.05 36.82 50.12 53.04 22.75 24.23
6 Performance Comparison
In this section, we present a detailed evaluation of world models for autonomous driving based on their performance across various tasks and metrics. Drawing from our previous discussions in Sec. 5, we benchmark representative algorithms to provide empirical evidence of their strengths and limitations.
6.1 Evaluation Platforms
NuScenes. Experiments are conducted on the widely used nuScenes [122] dataset with occupancy annotations provided by Occ3D [121]. This dataset contains 700 training sequences and 150 validation sequences, each with approximately 40 frames sampled at 2 Hz. The perception range is $[-40m,-40m,-1m,40m,40m,5.4m]$ and the voxel size is set to $[0.4m,0.4m,0.4m]$, leading to a grid size of $[200,200,16]$. Each grid cell is assigned one of 17 possible semantic categories, although some methods exclude ambiguous classes such as “other” and “other flat” during evaluation.
TABLE II: 4D occupancy forecasting performance on the Occ3D-nuScenes [121, 122] dataset (Sec. 6.3). Aux. Sup. denotes auxiliary supervision apart from the ego trajectory. Avg. denotes the average performance of that in 1s, 2s, and 3s.
  Method Input Aux. Sup. mIoU $\uparrow$ IoU $\uparrow$
1s 2s 3s Avg. 1s 2s 3s Avg.
Copy&Paste 3D-Occ None 14.91 10.54 8.52 11.33 24.47 19.77 17.31 20.52
OccWorld [47] 3D-Occ None 25.78 15.14 10.51 17.14 34.63 25.07 20.18 26.63
RenderWorld [106] 3D-Occ None 28.69 18.89 14.83 20.80 37.74 28.41 24.08 30.08
OccLLaMA-O [126] 3D-Occ None 25.05 19.49 15.26 19.93 34.56 28.53 24.41 29.17
DOME-O [117] 3D-Occ None 35.11 25.89 20.29 27.10 43.99 35.36 29.74 36.36
DFIT-OccWorld-O [50] 3D-Occ None 31.68 21.29 15.18 22.71 40.28 31.24 25.29 32.27
[…]
TPVFormer [123]+SelfOcc [132]+OccWorld-S [47] Camera None 0.28 0.26 0.24 0.26 5.05 5.01 4.95 5.00
OccWorld-F [126] Camera None 8.03 6.91 3.54 6.16 23.62 18.13 15.22 18.99
OccLLaMA-F [126] Camera None 10.34 8.66 6.98 8.66 25.81 23.19 19.97 22.99
RenderWorld [106] Camera None 2.83 2.55 2.37 2.58 14.61 13.61 12.98 13.73
DOME-F [117] Camera None 24.12 17.41 13.24 18.25 35.18 27.90 23.435 28.84
DFIT-OccWorld [50] Camera 3D-Occ 13.38 10.16 7.96 10.50 19.18 16.85 15.02 17.02
6.2 Perception in Static Scenes: 3D semantic occupancy prediction
The primary objective of this research is to accurately perceive and represent static elements in 3D space, forming the foundation for essential autonomous driving tasks such as scene segmentation, path planning, and navigation. By enabling precise and reliable reconstructions of static environments, this work supports critical functionalities of autonomous systems, including efficient route optimization, obstacle avoidance, and environmental understanding. These capabilities are pivotal for downstream processes, ensuring robust pathfinding and comprehensive environmental analysis in varied operational scenarios.
Metrics. Intersection-over-Union (IoU) and mean IoU (mIoU) are metrics for 3D semantic occupancy reconstruction and prediction. Higher IoU/mIoU values indicate more accurate capture of 3D geometry and semantics.
Results. As shown in Table I, the analysis highlights the performance of various methods on 3D semantic occupancy prediction tasks. Methods trained with 3D occupancy ground truths, (e.g., DOME [117]), achieve state-of-the-art performance (mIoU around 83.1%). Others like OccLLAMA [126] and OccWorld [47] report mIoU values of roughly 75.2% and 65.7%, respectively. In contrast, 2D-based methods [106, 127] generally reach lower mIoU ranges (20–30%); however, RenderWorld [106] still attains a competitive 27.87%, surpassing other 2D methods. Certain models excel in segmenting vehicles (cars, trucks, etc.) and environmental classes (sidewalk, vegetation) while encountering more difficulties with small objects such as bicycles and pedestrians.
TABLE III: Motion planning performance on the nuScenes [122] dataset (Sec. 6.4). Aux.Sup. denotes auxiliary supervision apart from the ego trajectory.
  L2 (m) $\downarrow$ Collision Rate (%) $\downarrow$
Method Input Aux. Sup. 1s 2s 3s Avg. 1s 2s 3s Avg.
IL [133] LiDAR None 0.44 1.15 2.47 1.35 0.08 0.27 1.95 0.77
NMP [64] LiDAR Box & Motion 0.53 1.25 2.67 1.48 0.04 0.12 0.87 0.34
FF [61] LiDAR Freespace 0.55 1.20 2.54 1.43 0.06 0.17 1.07 0.43
EO [105] LiDAR Freespace 0.67 1.36 2.78 1.60 0.04 0.09 0.88 0.33
[…]
UniAD [134] Camera Map & Box & Motion & Tracklets & Occ 0.48 0.96 1.65 1.03 0.05 0.17 0.71 0.31
UniAD+DriveWorld [18] Camera Map & Box & Motion & Tracklets & Occ 0.34 0.67 1.07 0.69 0.04 0.12 0.41 0.19
VAD-Tiny [135] Camera Map & Box & Motion 0.60 1.23 2.06 1.30 0.31 0.53 1.33 0.72
VAD-Base [135] Camera Map & Box & Motion 0.54 1.15 1.98 1.22 0.04 0.39 1.17 0.53
DriveDreamer [11] Camera Map & Box & Motion - - - 0.29 - - - 0.15
GenAD [19] Camera Map & Box & Motion 0.36 0.83 1.55 0.91 0.06 0.23 1.00 0.43
OccNet [136] Camera 3D-Occ & Map & Box 1.29 2.13 2.99 2.14 0.21 0.59 1.37 0.72
OccWorld-T [47] Camera Semantic LiDAR 0.54 1.36 2.66 1.52 0.12 0.40 1.59 0.70
OccWorld-S [47] Camera None 0.67 1.69 3.13 1.83 0.19 1.28 4.59 2.02
LAW [109] Camera None 0.26 0.57 1.01 0.61 0.14 0.21 0.54 0.30
Drive-OccWorld [49] Camera None 0.32 0.75 1.49 0.85 0.05 0.17 0.64 0.29
ViDAR [17] Camera None - - - 0.91 - - - 0.23
OccWorld-F [126] Camera Occ 0.45 1.33 2.25 1.34 0.08 0.42 1.71 0.73
OccLLaMA-F [126] Camera Occ 0.38 1.07 2.15 1.20 0.06 0.39 1.65 0.70
RenderWorld [106] Camera Occ 0.48 1.30 2.67 1.48 0.14 0.55 2.23 0.97
DFIT-OccWorld-V [50] Camera Occ 0.42 1.14 2.19 1.25 0.09 0.19 1.37 0.55
[…]
OccWorld [47] 3D-Occ None 0.43 1.08 1.99 1.17 0.07 0.38 1.35 0.60
RenderWorld [106] 3D-Occ None 0.35 0.91 1.84 1.03 0.05 0.40 1.39 0.61
OccLLaMA-O [126] 3D-Occ None 0.37 1.02 2.03 1.14 0.04 0.24 1.20 0.49
DFIT-OccWorld-O [50] 3D-Occ None 0.38 0.96 1.73 1.02 0.07 0.39 0.90 0.45
6.3 Perception in Dynamic Scenes: 4D Occupancy Forecasting
This subsection examines models’ capabilities to perceive and predict dynamic scenes, focusing on how moving objects and their interactions evolve over time. A central component of this task is 4D occupancy forecasting, which emphasizes temporal consistency for accurately capturing multi-agent dynamics. The aim is to predict future states of both moving objects and their interactions in complex traffic scenarios – a fundamental requirement for autonomous driving systems. By enabling reliable forecasting of dynamic environments, 4D occupancy models provide accurate predictions of object movements and interactions, forming the basis for crucial downstream tasks such as motion planning, collision avoidance, and safe navigation. These capabilities are essential in complex, multi-agent settings, ensuring that autonomous systems can adapt to rapidly changing conditions. Through maintaining temporal consistency and anticipating future scene dynamics, 4D occupancy forecasting also supports more robust decision-making and overall environmental understanding.
Metrics. In 4D occupancy forecasting, models predict future 3D occupancy from historical occupancy sequences to capture the scene’s evolution over time. Similar to Sec. 6.2, the evaluation relies on mIoU (mean Intersection over Union) and IoU (Intersection over Union) to gauge how accurately each future frame’s semantic occupancy is recovered, while placing additional emphasis on temporal accuracy and consistency across multiple time horizons (e.g., 1s, 2s, 3s).
Results. Table II summarizes the 4D occupancy forecasting performance on Occ3D-nuScenes, where predictions for 1s, 2s, and 3s into the future are assessed via mIoU and IoU. Notably, DOME-O attains state-of-the-art results (27.10% mIoU and 36.36% IoU), surpassing baseline methods such as OccWorld and RenderWorld by substantial margins. Even the purely camera-based DOME-F variant remains highly competitive, reflecting the model’s robustness in scenarios without direct 3D occupancy supervision.
6.4 Planning in Driving Scenarios: Motion planning
Motion planning is a critical component of autonomous driving, tasked with generating efficient, collision-free trajectories under real-time constraints. By accounting for both static and dynamic elements (e.g., obstacles, road geometry, and other vehicles) it enables safe navigation through complex environments including intersections, highway merges, and lane changes. Moreover, it supports energy-efficient routing strategies, thereby reducing fuel consumption or extending electric vehicle range. Effective motion planning not only underpins essential tasks like path generation and obstacle avoidance but also provides the foundation for broader applications, from warehouse robotics to urban delivery systems, where precise trajectory control ensures operational safety and efficiency.
Metrics. The evaluation of motion planning methods centers on key aspects such as route adherence, collision avoidance. Specifically, we adopt L2 error and collision rate as our core metrics: L2 error quantifies how closely a planned trajectory tracks the reference or desired path, while collision rate measures the frequency of unsafe interactions with obstacles. These metrics collectively ensure that the generated trajectories are both accurate and safe, supporting reliable navigation in dynamic driving scenarios.
Results. Table III presents a quantitative comparison of motion planning methods on the nuScenes [122] dataset, encompassing various sensor inputs (LiDAR, camera, and 3D occupancy) alongside different levels of auxiliary supervision (maps, bounding boxes, etc.). End-to-end autonomous driving frameworks (e.g., UniAD [134]) display strong results in both trajectory accuracy and collision avoidance, especially when trained on rich annotations (map, box, and motion supervision). By contrast, occupancy-driven methods (e.g., OccWorld [47], OccNet [136], RenderWorld [106], DFIT-OccWorld [50]) reduce reliance on auxiliary data, yet remain competitive in purely camera-based scenarios – indicating robust performance under more constrained conditions.
Furthermore, RenderWorld achieves approximately a 34% reduction in collision rate over a 3-second horizon compared to OccWorld, underscoring its capacity to forecast safer, long-term trajectories. Overall, the table highlights how both occupancy-centric and end-to-end solutions excel at generating precise, collision-free paths, with top-performing models striking an effective balance between minimal supervision and accurate motion forecasting.
7 Future Research Directions
With the rapid development of world models in the field of autonomous driving, the future research directions present a broad space for innovation. This chapter will focus on key frontier areas such as self-supervised learning, multi-modal fusion, advanced simulation, and efficient models, exploring how to further promote the development of autonomous driving systems in reducing dependence on labeled data, optimizing perception and decision-making, enhancing simulation realism, and achieving efficient deployment.
7.1 Self-Supervised World Models
Self-supervised learning (SSL) has demonstrated remarkable potential in reducing the dependency on annotated data, as evidenced by UniPAD [26]. Nevertheless, further research is necessary to not only minimize labeling costs but also to extract richer, domain-specific representations, as underscored by COPILOT4D [107] and SSR [108], which leverage discrete diffusion or sparse token-based representations to model complex spatio-temporal factors.
Reducing Label Dependency. Future efforts may explore coupling SSL with decoupled dynamic-flow mechanisms or voxel deformation approaches (e.g., DFIT-OccWorld [50]) to learn high-level scene dynamics without explicit ground-truth annotations, echoing the occupancy-centric strategies in UnO [29] and EO [105]. In addition, large-scale generative tasks such as video prediction (InfinityDrive [28]) or multimodal reconstruction (CarFormer [35]) can serve as powerful self-supervised objectives that capture complex 4D structures and reduce susceptibility to labeling errors – mirroring the rendering-based paradigm in RenderWorld [106] and the integrated BEV approach in BEVWorld [27].
Exploring Unlabeled Data Potential. Emerging methods like Think2Drive [99] and Symphony [137] illustrate how reinforcement learning (RL) agents and generative models can leverage vast unlabeled or partially labeled datasets to discover underlying spatiotemporal structures. By fusing simulation-based “thinking” with real-world sensor data, future frameworks could build on the latent forecasting strategies in LAW [109] or the temporal alignment techniques in ViDAR [17], thereby capturing subtle dynamics – from minor lane deviations to complex urban interactions – in a self-supervised manner. Such advances would not only reduce the cost of annotation but also sharpen domain-relevant representations, paving the way for safer and more robust autonomous driving systems.
7.2 Multi-Modal World Models
Multi-modal world models demand the integration of complementary sensor inputs (e.g., LiDAR, cameras, radar) to ensure robust perception and decision-making in real-world driving scenarios. Recent methods such as MuVO [48], UniPAD [26] underscore the benefits of fusing 2D and 3D streams, yet their reliance on high-capacity architectures and complex synchronization pipelines highlights the associated engineering challenges. Although RenderWorld [106] and ViDAR [17] demonstrate early success by merging camera, LiDAR, and BEV features, larger-scale integration into a single latent space (e.g., BEVWorld [27]) promises richer cross-modal cues and more holistic scene representations. CarFormer [35] leverage slot attention for object-centric tokenization across multimodal data, while UniPAD [26] combines volumetric differentiable rendering with LiDAR inputs for unified 2D–3D representation learning. Furthermore, large-scale transformer architectures such as DrivingGPT [100] and Token [138] push this paradigm by unifying image, point cloud, and map-based streams into a cohesive token space, thereby enabling simultaneous perception, prediction, and planning. Looking ahead, future research may extend these frameworks to incorporate emerging sensors (e.g., thermal or event cameras), aiming to enhance performance under adverse weather or low-visibility conditions. Additionally, leveraging large-scale multi-modal models to develop unified world models emerges as another promising direction, as demonstrated by approaches like DrivingGPT and Tokenize.
7.3 Advanced Simulation
Cross-Scenario Generalization. Autonomous driving systems require the capability to adapt across diverse road conditions, traffic densities, and cultural driving norms. To address this, DrivingDojo [114] provides richly annotated video clips for interactive world models, enabling more robust multi-agent interplay under varied traffic rules. Concurrently, OODGen [115] explores text-guided out-of-distribution scenario creation, pushing models to handle unpredictable conditions through synthetic data augmentation. Additionally, SimGen [113] merges real-world and simulator data to capture a wide range of phenomena, such as adverse weather and region-specific driving customs, thereby enhancing scenario adaptability. These efforts collectively aim to unify multi-agent interaction modeling, region-specific regulations, and adverse weather conditions within a single, domain-agnostic framework, ensuring consistent performance across heterogeneous traffic ecosystems. As generative approaches continue to evolve to accommodate broader environmental factors, future research is set to focus on integrating physical constraints, cultural factors, and dynamic interactions into more comprehensive, domain-agnostic pipelines.
Diffusion-based Generation. Recent advances in diffusion models drive the synthesis of high-fidelity and controllable driving scenes. OccSora [34] introduces 4D occupancy-based diffusion, enabling realistic and temporally consistent 3D simulations for enhanced decision-making. Similarly, Panacea [1] leverages panoramic diffusion to produce multi-view videos for training robust perception models, while InfinityDrive [28] breaks temporal constraints by extending scene generation to longer time horizons. Additional diffusion-based frameworks like DriveDreamer-2 [116] and DriveDreamer4D [25] improve data diversity by incorporating large language models or 4D reconstructions, respectively. Further, WORLDSIMBENCH [119] proposes a dual evaluation approach for video generation (assessing both perceptual quality and control-level realism), while WoVoGen [118] builds on a world volume-aware diffusion process to create multi-camera street-view videos with high spatio-temporal consistency. These methods collectively demonstrate the potency of diffusion-based architectures in synthesizing actionable, lifelike scenarios pivotal for autonomous driving research. By enabling robust control over scene realism and variability, diffusion-based generation paves the way for next-generation simulation pipelines that narrow the gap between virtual training environments and on-road performance.
Real-World Validation. Bridging advanced simulation with real-world performance remains a central challenge. Imagine-2-Drive [39] integrates a high-fidelity world model with a multi-modal diffusion policy actor for precise trajectory planning in simulated environments, showcasing how realism in simulations benefits downstream control tasks. Likewise, HoloDrive [37] unifies camera images and LiDAR point clouds to generate consistent 2D–3D datasets, closing the gap for perception tasks in real-world conditions. ReconDreamer [38] emphasizes online restoration techniques to maintain temporal coherence, and Vista [139] proposes high-resolution, versatile control for extended driving videos. Moreover, Delphi [40] focuses on generating controllable long-horizon sequences that promote stable decision-making in real-world deployments. By seamlessly combining these generative advances with physically realistic simulators (e.g., CARLA, LidarSim), researchers move closer to closed-loop systems that continuously refine perception, prediction, and control. Collectively, these developments pave the way toward robust, validated, and simulation-rich pipelines that bridge high-fidelity scene synthesis with reliable on-road performance.
7.4 Efficient World Models
UniPAD [26] and UniWorld [111] demonstrate how unified latent representations can streamline multi-task autonomous driving pipelines, covering detection, segmentation, and trajectory prediction in one framework. In parallel, large language model (LLM)-based approaches (e.g., DriveSim [101], DrivingGPT [100]) integrate textual and map-based information with sensor outputs, enabling a holistic view of the driving environment and potentially reducing the need for multiple specialized modules. Despite these advances, balancing model complexity with real-time constraints remains pivotal for large-scale, real-world deployments. DFIT-OccWorld [50] adopts decoupled flow and voxel deformation to model scene dynamics efficiently, underscoring a growing shift toward lightweight, expressive networks. Hierarchical scene representations such as Fiery [30] leverage multi-level feature maps to reduce computational overhead, while dynamic architectures like NeMo [46] adjust capacity on the fly based on input complexity. Going forward, integrating these efficiency-focused innovations within end-to-end driving pipelines promises to balance model complexity, latency, and reliability – key considerations for safely and scalably deploying autonomous vehicles.
8 Conclusion
World models have rapidly become a cornerstone for autonomous driving, enabling deeper integration among perception, prediction, and decision-making. Recent advances in multi-modal fusion unify data from cameras, LiDAR, and other sensors, while self-supervised learning and large-scale pretraining reduce dependence on annotated datasets. Generative methods, particularly diffusion-based approaches, now facilitate diverse synthetic data for long-tail scenarios, enhancing model robustness in rare or extreme conditions. New frameworks tightly couple motion prediction with planning algorithms, moving toward closed-loop paradigms that promise safer, more adaptive navigation. As sensing technologies evolve and cross-domain datasets proliferate, world models are poised to become even more integral to reliable, large-scale deployment of next-generation autonomous driving systems.

### v2

@@
 Abstract
-Recent breakthroughs in autonomous driving have revolutionized the way vehicles perceive and interact with their surroundings.
+Recent breakthroughs in autonomous driving have been propelled by advances in robust world modeling, fundamentally transforming how vehicles interpret dynamic scenes and execute safe decision-making.
 In particular, world models have emerged as a linchpin technology, offering high-fidelity representations of the driving environment that integrate multi-sensor data, semantic cues, and temporal dynamics.
-Such models unify perception, prediction, and planning, thereby enabling autonomous systems to make rapid, informed decisions under complex and often unpredictable conditions.
-Research trends span diverse areas, including 4D occupancy prediction and generative data synthesis, all of which bolster scene understanding and trajectory forecasting.
-Notably, recent works exploit large-scale pretraining and advanced self-supervised learning to scale up models’ capacity for rare-event simulation and real-time interaction.
-In addressing key challenges – ranging from domain adaptation and long-tail anomaly detection to multimodal fusion – these world models pave the way for more robust, reliable, and adaptable autonomous driving solutions.
-This survey systematically reviews the state of the art, categorizing techniques by their focus on future prediction, behavior planning, and the interaction between the two.
-We also identify potential directions for future research, emphasizing holistic integration, improved computational efficiency, and advanced simulation.
-Our comprehensive analysis underscores the transformative role of world models in driving next-generation autonomous systems toward safer and more equitable mobility.
+This paper systematically reviews recent advances in world models for autonomous driving, proposing a three-tiered taxonomy: 1) Generation of Future Physical World, covering image-, BEV-, OG-, and PC-based generation methods that enhance scene evolution modeling through diffusion models and 4D occupancy forecasting; 2) Behavior Planning for Intelligent Agents, combining rule-driven and learning-based paradigms with cost map optimization and reinforcement learning for trajectory generation in complex traffic conditions; 3) Interaction Between Prediction and Planning, achieving multi-agent collaborative decision-making through latent space diffusion and memory-augmented architectures.
+The study further analyzes training paradigms including self-supervised learning, multimodal pretraining, and generative data augmentation, while evaluating world models’ performance in scene understanding and motion prediction tasks.
+Future research must address key challenges in self-supervised representation learning, long-tail scenario generation, and multimodal fusion to advance the practical deployment of world models in complex urban environments.
+Overall, our comprehensive analysis provides a theoretical framework and technical roadmap for harnessing the transformative potential of world models in advancing safe and reliable autonomous driving solutions.
 Index Terms:
@@
 Equally pivotal is the system’s resilience in extreme or long-tail scenarios (e.g., severe weather, construction zones, or erratic driving behaviors), where performance shortfalls can compromise overall safety [10, 11].
-Within this context, constructing robust and stable world models has emerged as a foundational element.
+Within this context, constructing robust and stable world models [12] has emerged as a foundational element.
 The notion of a world model involves creating a high-fidelity representation of the driving environment – encompassing static structures (e.g., roads, buildings) and dynamic entities (e.g., vehicles, pedestrians) [3, 8].
@@
 Guided by the principle that the world model is central to the understanding of dynamic scenes, this survey aims to provide a comprehensive, structured review of existing methodologies.
-We categorize state-of-the-art research into three key areas: Future Prediction of the Physical World: Focusing on the physical world evolution of both dynamic objects and static entities [11, 25]; Behavior Planning for Intelligent Agents: Examining generative and rule-based planning methods that produce safe, efficient paths under uncertain driving conditions [12, 13]; Interaction Between Behavior Planning and Future Prediction: Highlighting how unified frameworks can capture agent interactions and leverage predictive insights for collaborative optimization [26, 27, 18].
+We categorize state-of-the-art research into three key areas: Generation of Future Physical World: Focusing on the physical world evolution of both dynamic objects and static entities [11, 26]; Behavior Planning for Intelligent Agents: Examining generative and rule-based planning methods that produce safe, efficient paths under uncertain driving conditions [13, 14]; Interaction Between Behavior Planning and Future Prediction: Highlighting how unified frameworks can capture agent interactions and leverage predictive insights for collaborative optimization [27, 28, 19].
 Specifically, we provide:
-  • An In-Depth Analysis of Future Prediction Models: We discuss how Image-/BEV-/OG-/PC-based methods achieve geometric and semantic fidelity in dynamic scenes, including 4D occupancy forecasting and diffusion-based generation.
+  • An In-Depth Analysis of Future Prediction Models: We discuss how Image-/BEV-/OG-/PC-based generation methods achieve geometric and semantic fidelity in dynamic scenes, including 4D occupancy forecasting and diffusion-based generation.
   • Investigation of Behavior Planning: We explore the behavior planning through both rule-based and learning-based approaches, demonstrating notable improvements in robustness and collision avoidance.
@@
 With the expanding research landscape and the urgency of real-world adoption, this survey aspires to serve as a valuable reference point for researchers and practitioners, laying the groundwork for safer, more robust autonomous driving solutions.
-§1§2§3§4§5§6§7§8§3.1§3.2§3.3§4
+§1§2§3§4§5§6§7§8§2§3.1§3.2§3.3§4§5
 Figure 1: Structure of the overall review.
-The first row shows the structure of the paper.
-The second and third rows detail future prediction of the physical world, behavior planning for intelligent agents, and interaction between behavior planning and future prediction.
-The fourth row highlights various methodologies for training models in autonomous driving, covering both self-supervised learning paradigms, pretraining strategies, and innovative approaches for data generation.
+The top row outlines the paper’s organization.
+The second and third rows illustrate the background and key components – generation of future physical world, behavior planning for intelligent agents, and the interaction between them.
+The fourth row highlights various methodologies for training models in autonomous driving, covering self-supervised learning paradigms, pretraining strategies, and innovative approaches for data generation.
+The bottom row showcases four application areas for world models in autonomous driving: scene understanding, motion prediction, simulation, and end-to-end driving.
 1.3 Structure
@@
 3 details the taxonomy of methods: Sec.
-3.1 delves into methods for future prediction of the physical world, discussing physical world evolution of dynamic objects and static entities.
+3.1 delves into methods for generation of future physical world, discussing physical world evolution of dynamic objects and static entities.
 Sec.
@@
 In this section, we first provide a detailed problem formulation for world models in autonomous driving (Sec.
-2.1), encompassing two key tasks: future prediction of the physical world and behavior planning for intelligent agents.
+2.1), encompassing two key tasks: generation of future physical world and behavior planning for intelligent agents.
 Then, in Sec.
@@
 $\!\!\!\!\bm{z}^{T+1},\!\tau^{T+1}\!=\!\bm{w}((\bm{I}^{T},\!\cdots,\!\bm{I}^{T-t}),\!(\bm{P}^{T},\!\cdots,\!\bm{P}^{T-t})).$ (1)
-The first core task is future prediction of the physical world [29, 27, 17, 30], which involves forecasting the future states of dynamic entities such as vehicles, pedestrians, and traffic elements.
+The first core task is generation of future physical world [30, 28, 18, 31], which involves forecasting the future states of dynamic entities such as vehicles, pedestrians, and traffic elements.
 This task emphasizes capturing potential interactions, stochastic behaviors, and uncertainties within rapidly changing and complex scenes.
@@
 Despite these advancements, balancing high-fidelity outputs with computational efficiency remains an active research challenge.
-2.2.3 Spatiotemporal Modeling Techniques
-Spatiotemporal modeling addresses how entities evolve in space and time within dynamic driving environments.
-Transformers. employ attention mechanisms to capture long-range temporal dependencies and multi-agent interactions, facilitating improved motion prediction and scene understanding [35].
-Slot Attention, on the other hand, segments the environment into interpretable entities, enabling more structured representations for downstream tasks such as behavior planning and motion forecasting [11].
-These approaches have proven valuable in addressing dynamic and complex driving scenarios.
-However, developing models that are both computationally efficient and robust in real-time settings remains a core challenge.
-This section provides a foundational understanding of world models in autonomous driving, highlighting how representation spaces, generative methods, and advanced modeling techniques together address the critical issues of uncertainty, complexity, and real-time operation.
 3 Taxonomy
-In this section, we classify the methodologies presented in the uploaded works into three primary categories, highlighting their unique contributions and interconnections: (1) Future Prediction of the Physical World, (2) Behavior Planning for Intelligent Agents, and (3) Interaction Between Behavior Planning and Future Prediction.
+In this section, we classify the methodologies presented in the uploaded works into three primary categories, highlighting their unique contributions and interconnections: (1) Generation of Future Physical World, (2) Behavior Planning for Intelligent Agents, and (3) Interaction Between Behavior Planning and Future Prediction.
 This section provides a structured understanding of the diverse approaches in autonomous driving research.
@@
 These frameworks collectively address the multifaceted challenges of autonomous navigation, driving innovation in safety, efficiency, and robustness.
-3.1 Future Prediction of the Physical World
-3.1.1 Image Representations
-Image representation-based approaches to future prediction in autonomous driving capitalize on advanced generative models to address challenges like limited data availability and complex scene dynamics.
-Techniques such as diffusion models and generative adversarial networks (GANs) enable the synthesis of high-fidelity images, enriching training datasets and enhancing the robustness of downstream perception and planning modules.
-For instance, DrivingDiffusion [14] deploys a diffusion model to generate temporally coherent video frames, creating diverse training scenarios that bolster object detection and motion segmentation in dynamic environments.
-By employing cascaded generation modules and localized prompts, DrivingDiffusion maintains spatial and temporal consistency, which is pivotal for reliably modeling traffic agents in motion.
-DriveDreamer [11] further integrates image synthesis and trajectory prediction to simulate realistic urban driving conditions, especially high-density intersections.
-This focus on interactive elements and unpredictable traffic flows enriches the planning process, enabling vehicles to anticipate and adapt to sudden changes on the road.
-Meanwhile, ViDAR [17] leverages GANs to synthesize future frames from monocular video sequences, employing temporal consistency layers to mitigate artifacts in moving objects and adapt to varying environmental factors such as weather changes.
-Building on these foundations, DrivingWorld [36] proposes a GPT-based video generation framework that unifies detection, prediction, and planning tasks within a single model, thereby streamlining data processing and decision-making.
-Complementarily, HoloDrive [37] synthesizes 2D–3D multi-modal scenes to reinforce the alignment between image and LiDAR data, enhancing predictive accuracy for more robust reconstruction of complex driving environments.
-ReconDreamer [38] focuses on the real-time reconstruction of dynamic scenes, emphasizing online restoration techniques that enable responsive modeling of evolving traffic conditions.
-In simulation environments, Imagine-2-Drive [39] advances high-fidelity world modeling in CARLA, equipping researchers and practitioners with more realistic test scenarios to evaluate autonomous systems under diverse conditions.
-Finally, Delphi [40] moves toward end-to-end autonomous driving by generating controllable, long-duration videos that facilitate broader generalization in complex traffic scenarios.
-Collectively, these image-based generative frameworks significantly strengthen the adaptability and safety of autonomous vehicles.
-They not only supply synthetic but highly realistic training data for perception tasks, but also refine predictive capabilities and planning strategies, enabling intelligent vehicles to navigate with greater awareness and robustness in rapidly shifting driving contexts.
-3.1.2 BEV Representations
+3.1 Generation of Future Physical World
+3.1.1 Image-based Generation
+Image representation-based approaches to future prediction in autonomous driving draw on advanced generative models, such as diffusion models, to address issues like limited labeled data and highly dynamic environments.
+By synthesizing high-fidelity images, these methods expand training datasets and bolster the resilience of downstream perception and planning modules.
+DriveDreamer [11] uses real-world driving data for controllable video generation in high-density urban traffic; DriveDreamer-2 [36] incorporates LLM-based prompts for customized multi-view video generation; DriveDreamer4D [26] leverages world model priors to produce spatial-temporally consistent 4D driving videos; ReconDreamer [37] integrates online restoration for accurate video reconstruction of dynamic scenes; WorldDreamer [38] predicts masked tokens to enable text-to-video and action-to-video generation; CarDreamer [39] provides an open-source platform supporting multi-modal video generation for autonomous driving.
+DrivingDiffusion [15] employs a latent diffusion model for layout-guided, multi-view driving scene video generation; Delphi [40] applies a controllable, long-duration diffusion-based approach to boost planning in autonomous driving videos.
+DrivingWorld [41] introduces a GPT-based framework for extended, high-fidelity driving video generation; GAIA-1 [16] uses unsupervised sequence modeling that integrates video, text, and action inputs for contextual driving video generation; DRIVESIM [42] leverages multimodal large language models to achieve robust, causal driving video predictions.
+HoloDrive [43] fuses camera and LiDAR data for future-frame driving video generation; BevWorld [28] applies cross-modal BEV learning for dynamic scene video generation; BEVGen [44] synthesizes street-level videos from semantic BEV layouts; Drive-WM[45] unifies visual forecasting and planning for consistent driving video generation.
+DriveArena [46] creates a closed-loop simulation for generative driving video; DrivingSphere [47] builds a 4D dynamic environment for multi-view driving video; Imagine-2-Drive [48] merges a high-fidelity world model with a diffusion-based policy network for diverse trajectory video generation; Vista [49] programmatically renders complex driving scenarios as high-quality videos; SimGen [50] conditions on simulators to produce varied driving videos for real-world generalization.LAW [51] trains a latent world model to predict future scene videos for end-to-end driving; Popov et al.
+[33] propose incremental latent data generation to combat environmental shifts in driving videos.
+Overall, these image-centric generative frameworks significantly enhance the adaptability and safety of autonomous vehicles.
+Beyond providing synthetic yet exceptionally realistic data for perception tasks, they also refine predictive capabilities and planning strategies, enabling intelligent vehicles to navigate rapidly changing road conditions with heightened awareness and robustness.
+3.1.2 BEV-based Generation
 Bird’s-Eye View (BEV) representations have emerged as a powerful paradigm for modeling future states in autonomous driving environments, thanks to their ability to unify multi-modal sensor data and facilitate structured, top-down scene understanding.
 As an alternative or complement to image-based approaches, BEV methods reduce redundant perceptual details from raw sensor readings, rendering high-level, spatially coherent views that are well-suited for motion prediction, occupancy modeling, and trajectory forecasting.
-A prominent illustration of BEV’s capabilities is GenAD [19], which introduces an end-to-end generative motion prediction pipeline via tokenized scene representations and a Variational Autoencoder (VAE).
-By capturing interactions between the ego vehicle and surrounding agents in a structured BEV space, GenAD yields accurate trajectory predictions even in cluttered traffic scenarios.
-Extending this representation to account for evolving semantic occupancy, BEVWorld [27] integrates a diffusion model to predict the geometric and semantic evolution of forthcoming scenes, merging multi-modal inputs (e.g., LiDAR and images) into a continuous latent space.
-In parallel, UNO [29] tackles 4D occupancy prediction from a purely unsupervised standpoint, leveraging historical signals to generate future point clouds and occupancy grids in a continuous fashion.
-Its implicit decoder mitigates quantization pitfalls and achieves state-of-the-art outcomes on benchmarks such as Argoverse and nuScenes.
-Similarly, FIERY [30] exploits spatiotemporal convolutional networks atop BEV inputs to model the dynamic evolution of traffic entities; by integrating probabilistic forecasting, FIERY can output a spectrum of possible future scenarios, which is particularly valuable for safety-critical tasks like instance segmentation and motion prediction.
-In addition to these representative works, a host of recent studies further highlights the diversity and adaptability of BEV-based world models.
-Methods like BevDepth [41] and PowerBEV [42] refine depth estimation and instance prediction within the BEV framework, whereas BevControl [43] and BEVGen [44] push the boundaries of multi-perspective consistency and generative street-view rendering, respectively.
-At the same time, InfinityDrive [28], GaussianWorld [45], and NeMo [46] expand upon volumetric, generative, and streaming occupancy strategies, illustrating the versatility of BEV-centric approaches when combined with next-generation neural architectures.
-Collectively, these advances underscore how BEV representations not only offer a unified canvas for perception, prediction, and planning but also significantly improve robustness in dynamic driving environments, paving the way for safer and more reliable autonomous systems.
-Generative scene synthesis has recently emerged as a valuable tool for enriching the data landscape used in behavior planning.
-For instance, BEVGen [44] synthesizes multi-view images from BEV segmentation masks, thereby enhancing data augmentation for tasks such as 3D detection and tracking.
-By manipulating the bird’s-eye view layout, BEVGen offers a structured way to generate diverse driving scenarios – helping learning-based planners acquire greater robustness and adaptability.
-Similarly, BEVControl [43] focuses on simplifying scene editing while maintaining geometric and visual consistency in the top-down view.
-This tool provides intuitive visual guidance for planning by allowing practitioners to modify complex elements of the scene with minimal effort, thus accelerating the iterative design and testing of learned planning strategies.
-Generative scene synthesis also contributes to behavior planning by enriching the data and visual references available.
-BEVGen [44] synthesizes multi-view images from BEV segmentation masks, enhancing data augmentation for tasks like 3D detection and tracking.
-BEVControl [43] simplifies scene editing while ensuring geometric and visual consistency, providing intuitive visual guidance for planning.
-These learning-based approaches underscore the transformative potential of data-driven behavior planning frameworks in addressing the complexities of autonomous driving, enabling safer and more adaptive decision-making in dynamic environments.
-3.1.3 OG Representations
+CarFormer [52] incorporates an object-centric BEV representation, splitting vehicles, pedestrians, and other entities into independent slots for more precise trajectory analysis.
+Building on similar structured representations, MILE [32] integrates a world model into imitation learning for multi-modal BEV predictions, while Popov et al.
+[33] further extend this pipeline by introducing latent-space generative models to mitigate covariate shift, ensuring semantic consistency between BEV and perspective views for enhanced robustness.
+GenAD [20] reframes autonomous driving as a generative modeling process, encoding scene elements in a BEV domain and sampling future trajectories through structured latent spaces.
+In parallel, UNO [30] focuses on unsupervised 4D occupancy field learning in BEV, enabling seamless adaptation to point cloud and BEV semantic occupancy forecasting, whereas FIERY [31] leverages spatiotemporal convolutions atop BEV inputs to predict multi-modal future instance trajectories under probabilistic uncertainty.
+PowerBEV [53] targets efficiency and stability by employing parallel multi-scale modules and flow-guided post-processing to reduce redundancy in BEV forecasting, achieving state-of-the-art performance on benchmarks like NuScenes with minimal model overhead.
+In contrast, BEVControl [54] addresses multi-perspective consistency by editing a BEV sketch layout, ensuring coherent geometry from top-down to perspective views—thereby offering an intuitive, flexible tool for scene manipulation and planning validation without directly outputting future trajectories.
+Aside from the aforementioned methods, to achieve substantial improvement in BEV semantic occupancy forecasting, there is another line of research that attempts to introduce a pre-text task called visual point cloud forecasting.
+ViDAR [18] proposes a novel pre-training framework for autonomous driving through a task termed visual point cloud forecasting—predicting future 3D point cloud sequences from historical visual inputs.
+With the aid of ViDAR, performance improvements are achieved in BEV generation.
+3.1.3 OG-based Generation
 Occupancy grid (OG) representations have become a cornerstone for predicting future states in autonomous driving, primarily due to their capacity to encode high-fidelity 3D spatial information.
@@
 However, such high-resolution modeling often incurs substantial memory and computational costs, especially in large-scale or real-time scenarios.
-The potential for sparse sensor data or irregular updates further complicates prediction accuracy, and controlled scenario generation typically relies on carefully curated trajectory conditions.
-A foundational approach in this domain is OccWorld [47], which introduces 4D occupancy grids to jointly predict semantic occupancy and agent trajectories in cluttered driving environments.
-Building on this groundwork, MUVO [48] integrates multi-modal data (i.e., LiDAR point clouds and camera streams) into high-resolution 3D voxel grids, elevating prediction quality in both vision-based and point cloud tasks.
-Extending temporal modeling, OccSora [34] leverages diffusion models to generate future occupancy states, enabling controlled scenario simulation for motion-planning pipelines.
-Further integrating prediction with decision-making, Drive-OccWorld [49] fuses occupancy forecasts with a motion planner, leveraging historical grid states and vehicle trajectories to optimize driving strategies in real-world contexts.
-Recent works push OG-based methods toward greater efficiency and scalability.
-DFIT-OccWorld [50] employs a decoupled dynamic flow mechanism and image-based supervision, achieving end-to-end, non-autoregressive occupancy forecasting while mitigating computational overhead.
-GaussianWorld [45] reframes OG prediction in a 4D Gaussian space, explicitly modeling scene evolution to enhance streaming performance on datasets like nuScenes without substantial resource demands.
-Across these advances, learning-based OG techniques consistently prove their worth in motion forecasting, scene reconstruction, and safe route planning.
-By balancing high-resolution occupancy outputs with emerging strategies for real-time efficiency, the latest research points toward solutions that can seamlessly adapt to the complexity and dynamism of real-world autonomous driving.
-3.1.4 PC Representations
+OG-based Generation (i.e., occupancy forecasting) originates from forecasting semantic occupancy grids on bird’s-eye view (BEV) [55, 56, 57, 58, 59, 60].
+It is employed to predict how the surrounding occupancy will evolve in the near future beyond the current moment.
+Then gradually expand to predict 4D occupancy grids.
+For instance, Occ4cast [61] aims to forecast a sequence of dense and completed occupancy grids under the Eulerian specification, offering a more comprehensive perception in complex dynamic environments.
+Recent OG-based Generation approaches begin to closely align with world models, showing a trend toward integrating multi-view images, language, and actions to achieve advanced 4D occupancy forecasting and planning.
+They adopt multi-modal sensors, generative architectures, and spatiotemporal tokenization strategies to comprehensively model future occupancy for decision-making in autonomous driving.
+Specifically, MUVO [62] pioneers multi-modal sensor fusion (camera + LiDAR) for actionable 3D occupancy prediction, while Cam4DOcc [63] establishes the first camera-only 4D occupancy benchmark with standardized protocols.
+Building on these, OccWorld [64] introduces GPT-style autoregressive generation for scene token evolution, and DFIT-OccWorld [65] streamlines training via decoupled dynamic-static voxel warping.
+Concurrently, DOME [66] adopts diffusion transformers for controllable long-horizon forecasting, whereas DriveWorld [19] and Drive-OccWorld [67] integrate memory-augmented world models with BEV-based planning, bridging occupancy forecasting and trajectory optimization.
+GaussianWorld [68] reformulates scene dynamics in 3D Gaussian space, enabling ego-motion-aligned prediction, while OccLLaMA [69] unifies vision-language-action modalities via occupancy tokenization.
+OccSora [35] extends this with diffusion-based 4D generation conditioned on trajectories, and RenderWorld [70] achieves vision-only efficiency via Gaussian Splatting and disentangled VAE encoding, culminating in a cost-effective, end-to-end autonomy pipeline.
+3.1.4 PC-based Generation
 3D point cloud (PC) representations, typically obtained from LiDAR sensors, have emerged as a linchpin in autonomous driving for tasks such as occupancy forecasting, dynamic scene modeling, and predictive reconstruction.
@@
 However, the sparsity and irregular sampling of LiDAR scans – combined with real-time computational constraints – continue to pose significant algorithmic challenges.
-Early generative efforts targeted realistic LiDAR data simulation for data augmentation and testing.
-For instance, lidarGeneration [51] adopted a GAN-based framework to mimic real-world LiDAR statistics, while Lidarsim [7] introduced LidarSim to construct near-realistic LiDAR scenes by leveraging empirical sensor distributions.
-Subsequent works further advanced the diversity and efficiency of point cloud generation.
-Lidargen [52] proposed a deep learning pipeline that preserves physical consistency while increasing point cloud variability.
-Meanwhile, AXform [53] employed attention mechanisms to map latent features into 3D space, and PSF [54] developed a straight-flow approach that accelerates the generation process without compromising fidelity.
-In parallel, NFL [55] introduced Neural LiDAR Fields for synthesizing novel-view point clouds, demonstrating effectiveness under dynamic or partially observed scenes.
-Extending neural rendering concepts, Nerf-lidar [56] proposed NeRF-LiDAR, leveraging neural radiance fields to produce high-resolution LiDAR data suitable for realistic 3D reconstructions.
-A foundational step toward future prediction with PC data is offered by 4DOcc [57], who recast point cloud forecasting as a 4D occupancy prediction problem.
-Rather than directly inferring raw point coordinates, they map sequential LiDAR scans into intermediate occupancy grids and employ temporal models (e.g., Transformers) to estimate future occupancy states.
-This paradigm neatly bypasses heavy reliance on meticulously annotated LiDAR data by rendering pseudo LiDAR sequences from the predicted occupancy volumes.
-Recent work further demonstrates the synergy of point cloud representations with advanced neural volumetric techniques.
-For instance, NeMo [46] leverage self-supervised 3D reconstruction and motion cues to achieve robust geometric and dynamic understanding – often outperforming 2D or BEV-based approaches in scenarios with frequent occlusions or varied terrains.
-Similarly, ViDAR [17] extends this idea by predicting future point cloud distributions from purely visual inputs, thus unifying semantic, spatial, and temporal learning.
-By integrating generative modeling with latent-space transformations, ViDAR demonstrates notable gains in 3D detection, motion forecasting, and route planning.
+PC-based Generation predicts future point clouds from past point cloud inputs, and existing methods [71, 72] directly forecast future laser points despite challenges posed by sparse and irregular LiDAR data.
+PointRNN [71] employs a recurrent neural network to capture spatiotemporal features in moving point clouds by incorporating point-level operations, while MoNet [72] leverages motion information to model dynamic variations in continuous point cloud sequences through embedding motion features into its network architecture.
+Another line of methods uses range images [73, 74], a representation obtained by projecting point clouds to dense 2D images using sensor intrinsic and extrinsic parameters.
+Based on historical range images, they apply 3D convolutions [75, 76] or LSTMs [77, 78] to predict future point clouds, yet they additionally model the motion of sensor intrinsic and extrinsic parameters.
+On the one hand, PC-based Generation have gradually evolved to target realistic LiDAR data simulation for data augmentation and testing.
+SE3-Nets [79] predicts a scene point cloud in the next frame by considering the action applied to an object in the scene, while lidarGeneration [80] uses VAEs and GANs to generate a scene point cloud from a random noise vector, and Dscnet [81] generates a high-resolution scene point cloud conditioned on stereo images and low-resolution LiDAR inputs.
+lidarGeneration [80] adopted a GAN-based framework to mimic real-world LiDAR statistics, whereas Lidarsim [7] constructs near-realistic LiDAR scenes by leveraging empirical sensor distributions.
+Subsequent works have further advanced the diversity and efficiency of point cloud generation: Lidargen [82] proposed a deep learning pipeline that preserves physical consistency while increasing point cloud variability, AXform [83] employed attention mechanisms to map latent features into 3D space, and PSF [84] developed a straight-flow approach that accelerates the generation process without compromising fidelity.
+In parallel, NFL [85] introduced Neural LiDAR Fields for synthesizing novel-view point clouds, demonstrating effectiveness under dynamic or partially observed scenes, and extending neural rendering concepts, Nerf-lidar [86] proposed NeRF-LiDAR, which leverages neural radiance fields to produce high-resolution LiDAR data suitable for realistic 3D reconstructions.
+On the other hand, Recent advancements in point cloud generation have increasingly intertwined with world models for autonomous driving.
+4DOcc [87] transforms sequential LiDAR scans into intermediate occupancy grids and utilizes temporal models, such as Transformers, to predict future occupancy states, thereby reducing dependence on extensively annotated LiDAR data.
+Copilot4D [88] tokenizes LiDAR observations using a Vector Quantized Variational Autoencoder (VQVAE) and forecasts future point clouds through discrete diffusion, enabling unsupervised learning of world models.
+NeMo [89] employs self-supervised 3D reconstruction and motion cues to achieve a robust understanding of geometry and dynamics, often surpassing 2D or Bird’s-Eye View (BEV)-based methods in complex environments.
+ViDAR [18] introduces a pre-training framework for autonomous driving by predicting future 3D point cloud sequences from historical visual inputs, effectively integrating semantic, spatial, and temporal learning.
 Collectively, these methodologies highlight the effectiveness of 3D point cloud representations in autonomous driving.
@@
 The following explores learning-based planning, highlighting key contributions.
-A central concern in learning-based approaches is balancing performance with safety.
-SafeDreamer [98] integrates Lagrangian methods into the Dreamer framework to form a safe reinforcement learning algorithm, achieving near-zero safety violations across multiple tasks in the Safety-Gymnasium benchmark.
-This highlights a growing trend toward incorporating explicit safety constraints into RL-based planners.
-Another reinforcement learning–based approach, Think2Drive [99], introduces a neural planner capable of handling quasi-realistic scenarios using techniques such as the reset technique, automated scenario generation, termination-priority replay strategy, and steering cost function.
-This approach demonstrates that an agent can “think” through imagined scenarios in a latent world model to enhance learning efficiency.
-Tested in CARLA-v2, Think2Drive achieves higher sample efficiency and improved decision quality compared to standard RL baselines.
+Model-based reinforcement learning (MBRL) leverages world models to predict environmental states, improving data efficiency and making it suitable for autonomous driving due to high sample efficiency and learnable state transitions [128].
+Early attempts like MILE [32] introduced model-based approaches to autonomous driving, though their reliance on expert-collected data limited performance [128].
+MBOP [129] comprehensively addresses planning challenges by utilizing multiple offline-learned models within a model-predictive control (MPC) framework, enabling flexible reward function extensions and state constraint incorporation.
+While enhancing interpretability through learned dynamics models, MBOP employs simplistic deterministic models that ignore environmental stochasticity and aleatoric uncertainty [128].
+UMBRELLA [130] advances this paradigm by employing stochastic dynamics models to capture diverse traffic scene evolutions, explicitly addressing both epistemic and aleatoric uncertainties while learning from offline data.
+This approach integrates partial observability considerations and uses interpretable representations to tackle simultaneous prediction, planning, and control challenges in self-driving vehicles (SDVs).
+SafeDreamer [131] enhances safety considerations by integrating Lagrangian methods into the Dreamer framework, achieving near-zero safety violations in Safety-Gymnasium benchmarks.
+This reflects the growing emphasis on embedding explicit safety constraints into reinforcement learning (RL)-based planners for real-world applications.
+Think2Drive [128] models environmental transitions through world models [12], employing them as neural network simulators for latent-space planning.
+The method introduces a neural planner with reset techniques, automated scenario generation, and steering cost functions, demonstrating how agents can improve learning efficiency by ”thinking” through imagined scenarios in latent world models.
 A parallel thread focuses on large language models.
@@
 Finally, BEVWorld [28] unifies diverse sensor modalities within a single BEV latent space using diffusion modeling, showcasing the current push toward holistic, self-supervised world models that fuse complementary signals for perception, prediction, and planning in autonomous driving.
+AD-L-JEPA [137] introduces a self-supervised framework for autonomous driving using LiDAR data, leveraging a Joint Embedding Predictive Architecture (JEPA) to learn spatial world models by predicting BEV embeddings, eliminating generative/contrastive mechanisms and explicit data reconstruction while capturing occluded or uncertain environmental details.
+AD-L-JEPA achieves 5×5$\times$5 × faster pre-training than SOTA approaches (e.g., Occupancy-MAE [138], ALSO [139]) by avoiding contrastive pair curation and generative overhead, while demonstrating robust transfer learning even with partially randomized encoder initialization and superior label efficiency on downstream tasks like 3D object detection.
 4.2 Pretraining Strategies
@@
 Scene understanding serves as the foundation for real-time environmental perception and comprehension in autonomous driving.
-This technology enables autonomous systems to identify roads, vehicles, pedestrians, and traffic signs, thereby facilitating safe and effective driving decisions.
+This technology enables autonomous systems to identify roads, vehicles, pedestrians, and traffic signs—thereby facilitating safe and effective driving decisions.
 With the advancement of the field, world models are introduced to enhance the depth and breadth of scene understanding.
@@
 However, traditional simulation methods often fail to capture the full complexity of dynamic driving conditions, prompting researchers to explore world models – comprehensive frameworks that encompass both spatial and temporal aspects of the driving context.
-For instance, the OG Representation-based OccSora [34] is a diffusion-based 4D occupancy generation model designed to simulate the evolution of 3D scenes for autonomous driving.
+For instance, modern platforms like CARLA [148] (Unreal Engine-based) and AirSim [149] (Microsoft’s aerial/ground simulator) leverage game engines to synthesize high-fidelity environments with customizable sensors (LiDAR, cameras), dynamic weather, and traffic scenarios.
+While these tools enable physics-based interactions and modular testing, their reliance on synthetic data introduces a reality gap: overly perfect textures/lighting and scripted traffic patterns fail to capture natural imperfections or rare edge cases.
+Recently, another line of simulations, which are based on world models, is gaining ground.
+The OG Representation-based OccSora [35], is a diffusion-based 4D occupancy generation model designed to simulate the evolution of 3D scenes for autonomous driving.
 OccSora employs a 4D scene tokenizer to extract compact spatio-temporal representations, thereby enabling high-fidelity reconstruction of extended occupancy videos.
@@
 This trajectory-aware 4D generation serves as a world simulator to inform decision-making in autonomous vehicles.
-Another method, image Representation-based SimGen [113] learns to produce diverse driving scenes by integrating data from both simulators and the real world.
+In contrast, image Representation-based SimGen [50] learns to produce diverse driving scenes by integrating data from both simulators and the real world.
 Through a cascade diffusion pipeline, SimGen bridges the gap between simulated and real-world data by adhering to simulator-driven layout guidance and rich text prompts, ultimately yielding realistic driving scenarios.
@@
 End-to-end autonomous driving frameworks (e.g., UniAD [162]) display strong results in both trajectory accuracy and collision avoidance, especially when trained on rich annotations (map, box, and motion supervision).
-By contrast, occupancy-driven methods (e.g., OccWorld [47], OccNet [136], RenderWorld [106], DFIT-OccWorld [50]) reduce reliance on auxiliary data, yet remain competitive in purely camera-based scenarios – indicating robust performance under more constrained conditions.
+By contrast, occupancy-driven methods (e.g., OccWorld [64], OccNet [164], RenderWorld [70], DFIT-OccWorld [65]) reduce reliance on auxiliary data, yet remain competitive in purely camera-based scenarios—indicating robust performance under more constrained conditions.
 Furthermore, RenderWorld achieves approximately a 34% reduction in collision rate over a 3-second horizon compared to OccWorld, underscoring its capacity to forecast safer, long-term trajectories.
@@
 Future efforts may explore coupling SSL with decoupled dynamic-flow mechanisms or voxel deformation approaches (e.g., DFIT-OccWorld [65]) to learn high-level scene dynamics without explicit ground-truth annotations, echoing the occupancy-centric strategies in UnO [30] and EO [60].
-In addition, large-scale generative tasks such as video prediction (InfinityDrive [28]) or multimodal reconstruction (CarFormer [35]) can serve as powerful self-supervised objectives that capture complex 4D structures and reduce susceptibility to labeling errors – mirroring the rendering-based paradigm in RenderWorld [106] and the integrated BEV approach in BEVWorld [27].
+In addition, large-scale generative tasks such as video prediction (InfinityDrive [29]) or multimodal reconstruction (CarFormer [52]) can serve as powerful self-supervised objectives that capture complex 4D structures and reduce susceptibility to labeling errors—mirroring the rendering-based paradigm in RenderWorld [70] and the integrated BEV approach in BEVWorld [28].
 Exploring Unlabeled Data Potential.
 Emerging methods like Think2Drive [128] and Symphony [165] illustrate how reinforcement learning (RL) agents and generative models can leverage vast unlabeled or partially labeled datasets to discover underlying spatiotemporal structures.
-By fusing simulation-based “thinking” with real-world sensor data, future frameworks could build on the latent forecasting strategies in LAW [109] or the temporal alignment techniques in ViDAR [17], thereby capturing subtle dynamics – from minor lane deviations to complex urban interactions – in a self-supervised manner.
+By fusing simulation-based “thinking” with real-world sensor data, future frameworks could build on the latent forecasting strategies in LAW [51] or the temporal alignment techniques in ViDAR [18], thereby capturing subtle dynamics—from minor lane deviations to complex urban interactions—in a self-supervised manner.
 Such advances would not only reduce the cost of annotation but also sharpen domain-relevant representations, paving the way for safer and more robust autonomous driving systems.